A Complete Bibliography of Publications in IEEE Distributed Systems Online

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 1.12

Title word cross-reference

3 [App02d].

.NET [Bon06, Kar04b, LL03, Sto02e].

2 [BCA^{+}01]. 2.0 [Got08f, Mil07c]. 2000} [Abr03b]. 21st [Kor08, Mur01]. 2nd [Feu02, Sem02b, Zen02, CM06].

3rd [Dwi02b, Puz02b].

4th [She02a].

5.0 [Veg02a].

95 [Vil02].

= [CKB06].

Abbas [Puz03b]. Access [McL03a, Opp03, VPW^{+}06]. Accessing [RdCVM07]. Acclaimed [Wea01b]. Achilles [Sat03]. Act [Dav04b]. Action [Gon02]. Ad [FCBG07, Got02e, SGP04]. Adaptable [HMJV08]. Adaptation [PJMW08]. Adapting [PKP03]. Adaptive [DLBC04, FCBG07, FPA05, Gue04, KDH^{+}06, KRG08, MSCV05, OOS01a, WSKP01]. Add [Got02e]. Add-On [Got02e]. Address [Got06c]. Advanced [Sem03b]. Advantages [Lat05a]. Advertising [Cos03d, Sha03]. Advice [Mor03]. advocates [Alf04b]. After [Hed01a]. Agent [Gan07, GF08, GK02, KGR02, Kar04c]. Agent-Based [GK02]. Agents [Bad07, Gan06, Gan08, Huh03, PSB^{+}03, PSS02, Rod02, Xue01, Gue04]. AIDE [Wea01a]. Aims [Alf06b]. Airwaves [Got02a]. Alan [Mil07a, Puz03a]. Alesso [Vol02]. algorithm [ATS04]. Algorithmic [Rao08]. Algorithms [Par04, PBB02, Xue02b, Yeh06, Yeh07, Sas02a, Xue02b]. Alice [Sta02a]. Allen [Lin03a]. Already [Pie06]. Amazon.com [LSY03]. Analyzing [FH06]. Anderson [Sem02c, Vil03]. Andy [Sem02c]. Annotation [Paw06]. Answer [Sto02e]. Ant [Mor03, Mor03]. Appajodu [Bar03]. Applicability [Sto06a]. Application [Ber05b, Gal05, JRC^{+}06, Ram01, VPW^{+}06]. Application-Level [Ram01]. Application-Specific [VPW^{+}06]. Applications [AR07, Bar03, BCPK03, BJ00, Cha03, CL02, Dwi02b, Feu02, Fox03, FPA05, Kar02, KRG08, Kim03, MSCV05, MSC^{+}02, Mjø02, OSE06, PJMW08, Pet05, Pha01, SER^{+}04, SWH06, Sem03b, SBR^{+}07, SS06, She02b, Sta03b, Sta03a, TCB03, WSKP01, AvSQP05, HGG^{+}04b, HGG^{+}04a, KLZ04, Omm04, Cai03, Woo02]. Applied [Cha04a]. Applying [Sem06]. Approach [Pér03a, Rao08, SBR^{+}07, Xue02b]. Approaches [HM06]. Architectural [KRG08]. Architecture [GMG07, HAMC^{+}04, MSPB02, RCS08, Lin03a]. Architectures [DMJ^{+}02, Puz03b, She02a]. Area [ST07]. Argument [DFK08]. Art [Gan07, Mir02, OT07c, Cha03]. Artificial [Cha04b]. Arvind [Kar03]. ASN.1 [Ros03, Ros03]. ASP.NET2 [Zal04]. Aspects [Stp04]. Assertion [DSD07]. Assistant [DMJ^{+}02]. association [ATS04]. Attacks [Got07g, Cos04b]. Attention [Got06m]. Attitudes [ET03]. Attractive [Veg07]. Augment [Cos04d]. Augmented [DMJ^{+}02]. Augmented-Reality [DMJ^{+}02]. Authentic [OT06a]. Authentication [SK06]. Automated [AKS07]. Automatic [Rod02]. Autonomic [RK06, Toh04]. autonomic-computing [Toh04]. Autonomous [PJMW08]. Aviel [ET03]. Avison [Abr03a]. Aware [PMP03a, PMP03b, PMP03c, PMP03d, PMP03e, Sat02b, TCB03].

B [Mil00b, Puz03a]. B. [Hil03]. Back [App02a, GBK03, Sto02c]. Backbone [JG08]. Backlash [Got02d]. Balancing [Far02, OOS01a, OOS01b]. Bane [Opp03]. bans [Cos04a]. Bar [McG03d]. Barbara [Mil05]. Based [Gav01, Gee06, GK02, OT07e, OOS01b, PSB^{+}03, PLB07, PMF07, SWH06, Sch01c, SBR^{+}07, SSD^{+}04, Veg02b, Kar04c, Law04, Ano02b]. Basic [Ano02a]. Basics [Cai03]. Basis [Puz04]. Battery [Alf06b]. be [Cos03b]. Becomes [Got08f]. Beginner [SOB02, Yeh06, Yeh07]. Beginners [Dwi01, Pér03a]. Beginning [AB07]. Behavior [HWCG06]. Behind [Zal03]. Believe [Pie06]. Bell [HCW05]. Bellovin [ET03]. Bennett [She02b]. Best [Pet08b, Ros03]. Better [Sto02a]. Between [Ros03, TT03, SK06]. Beyond [Got06l, Hed00a, Kil02, Sti03, Got06b]. Bill [Dwi02a]. Bilson [Veg02a]. Bioinformatics [Lat07]. Biological [Cha04b]. Biology [Lat07]. Biometric [MM07]. Biometrics [Got08h, MM07]. Blake [Sem02a]. Blanket [Sto02g]. Blending [Gal05]. Blessing [Opp03]. Blew [Got06k]. Blinding [Got07c]. Blocks [Woo02]. Bluetooth [Sto01, Sto02c]. Blurring [SK06]. Bob [Abr03a]. Book [Abr03a, Abr03b, Abr06, Ano03, Bar03, Ber04, Bon03, Bon06, C.03, Cai03, Cha03, Cha04a, Dwi02a, Dwi02b, ET03, ET04b, ET06, Gan06, Gan07, Gor04b, Hil03, Iva03, Kac07, Kar02, Kar03, Kim03, Kow06, Lat06a, Lat06b, Law04, Lin02, Lin03a, Lin03b, Lin03c, Liu06, Mar08b, Mor03, Nec06, OT06a, OT06b, OT07a, OT07b, Pér03a, Pér03b, Puz02a, Puz02b, Puz03a, Puz03b, Puz03c, Puz04, Puž06a, Puz07a, Ros06, Ros03, Sed06, Sem02a, Sem03b, Sem03a, Sem03c, She02a, Sti03, Sto06b, Veg03, Vil03, Vol02, Vre04, Wil06, Woo02, Yan06, Zal03, vS01]. Boost [Alf06b, Got06d]. Border [Got08h]. Bottlenecks [Wil06]. bottom [Alf04a]. Boxer [Xue02b]. Breaking [Wil06]. Breakthrough [OT06a]. Breakthroughs [Got07a, Mil04]. Brian [Cha03]. Brick [Sto02a]. Brief [Alf06a, Cos02, Cos03a, Cos03c, Cos03b, Cos03d, Cos03e, Dav03, Gee06, McG03a, McG03b, McL02, McL06, Sha03, Sto02a, Sto02c, Sto02b, Sto02d, Sto03a, Sto03b, Vot03, Cor04]. Brilliant [Dwi02a]. brings [Cos04c]. Brink [Got07a]. Broadband [Lew02a, McL03a]. Broker [BAG01]. Brooks [Veg02a]. Brooks-Bilson [Veg02a]. Browsing [LHWX03]. Buege [Cha03]. Build [Pet08b]. Building [Kim03, RK06, Sem02c, Sto02a, Woo02, Omm04]. Bullies [McG03a]. Business [C.03, Dwi02a, Sem03b, Sto06a]. Buzz [McL03a]. Byrnes [C.03].

C [Vre04]. C# [LL03]. Caching [Vit01]. Calendar [PSS02]. Calishain [Veg03]. Call [vS00]. Camera [Cos04a]. Can [BM07]. Canning [Sto03a]. Capital [Got05]. Care [Sta02b]. Career [Mil05]. Carolina [WF06a, WF06b]. Case [BPBT06, FvdMAR08]. Catallactic [JRC^{+}06]. Catalyst [Sat02a]. CCM [WSKP01]. CDMA [Vil02, Vil02]. cdma2000 [Vil02]. Cell [Cos03e]. Cellular [Vil02]. Cellular/PCS [Vil02]. Center [Got06m]. Centered [CL02]. Centers [Meh02]. Centric [GGC03, Mid02]. Century [Kor08, Mur01]. Certification [Alf06b, Pér03b]. Chads [Fer01]. Challenge [CKB06]. Challenges [HM06, PMP03a, PMP03b, PMP03c, PMP03d, PMP03e, Sat02b, Var06]. Chandra [Stp02]. Change [Got08e, Hil03]. changes [Bak04]. Checkpoint [dCKC06]. Cheswick [ET03]. Chief [Sat03]. Ching [Got07h]. Chip [Got07a, SBR^{+}07]. Chris [She02b]. Christian [C.03]. Cisco [Puž06a]. Classical [Xue02a]. Classification [KLAR08]. Classroom [Alf05b, Wei00]. Claude [Abr03b]. Clayton [Pér03a]. Climate [SKA^{+}02]. Clock [BR02]. Cluster [BV07, SSD^{+}04]. Cluster-Based [SSD^{+}04]. Code [McG03d, Ano02a]. ColdFusion [Veg02a]. Collaboration [Epp03, SS06]. Collaborative [LSY03, Qin01, SER^{+}04, SS06, SYS07, Wea01a]. Collection [Lat07]. Colleges [Mur06]. Collide [Gan06]. Comer [Dwi02b, Puz02b, She02a]. Comes [Cos02]. Coming [Got06a, LCF^{+}08, Sem07]. Commerce [Che02, Gal05, Mjø02, Alf04a, Che02]. Communication [LRA01, Lin02, Ros03]. Communications [Lin02, Lin03b, Puz03a, Lin03b]. communities [RH04a]. Community [Got06b, Mur06, Pie06, PMF07, SS06, SK06, Bak04]. Companies [OT07e]. Comparison [HL07]. Compile [Paw06]. Compile-time [Paw06]. Compilers [Lin03a]. Complete [Sem03a, Sem03c]. Complex [Got08k, Iva03]. complexity [HMG05, vS04]. Component [Bon06, SSD^{+}04]. Component-Oriented [Bon06]. Components [Mir02]. Comprehensive [Dwi01, Pet08a]. Compulsive [PSB^{+}03]. computation [KLZ04, Mar05]. Computational [DPPR03, Gor04b, Lat07, LMH05, OT06c]. Computer [Abr06, App02a, BCPK03, DMJ^{+}02, Dwi02b, MSC^{+}02, Moo01, Pap06, Puz02b, STHK02, TCB03]. Computers [Dav04a, Hil03, Mir02, Ber04, BJ00, Mir02]. Computing [Abo03, AHC^{+}07, Alf06a, Ano01, AB07, App02c, BR02, Bir03, BAG01, BV07, BR07, Cai03, CFJ^{+}02, Che05b, CP03, Cla03, CKB06, DMJ^{+}02, DF03, Epp03, FH03, FvdMAR08, FHR^{+}08, FJH^{+}08, FBG^{+}08, Fox01, Fox02, Fox03, Fra03, Gen02, Got06e, Hed00a, Hed01a, HH03, HAMC^{+}04, Huh03, Jai02, Kar02, Kar04a, Kim02, Kor08, Kul07, Kum03, Lat05a, Lat05b, Lat07, LV03, LCF^{+}08, MSPB02, Meh02, Met03, Mil00b, Mir02, Mur06, Mur07, OT07c, PSB^{+}03, Pap06, PKP03, PN03, PMP03a, PMP03b, PMP03c, PMP03d, PMP03e, PHM03, Pra01, RK06, San02, San03, Sas02b, Sat02a, Sat03, SS06, SYS07, Sta02a, SKA^{+}02, Sta02b, Sta03b, Sta03a, Sto06a, Stp04, TT03, Vin02, Vin03c, Vin03b, Vin03a, WF06a, WF06b, Yan05, Yan07, Mar05, Mil04, Toh04, YPSW04, Kar02, Cha04a]. Computing [Sta02a]. Concept [Ano02a]. Concepts [ET03, Lat06a, Xue02a]. Concise [Pér03a]. Concurrency [Att07]. Concurrent [Lat06b, Par04]. Condor [AB07]. Configuration [Puž06a]. Conflict [Xue01]. Congress [McL02]. Congressional [Dav03]. Connection [Ram01]. Connections [DFK08, Pow01]. Connectivity [OT07a]. Consistent [MSPB02]. Console [Got06l]. Constraints [KRG08]. Container [RSL06]. Content [BPBT06, PKP03, Sto02h]. Context [RCS08, Sat02b, dRE06]. Context-Aware [Sat02b]. Context-Processing [RCS08]. Continuity [BGSK04]. continuous [HGG^{+}04b, HGG^{+}04a]. continuous-media [HGG^{+}04b]. Control [Bad07, BGSK04, PJMW08, VPW^{+}06]. Converge [McL06]. Converged [Puz06b]. Cookbook} [Bon03]. Cooking [Bon03]. Coordination [SWH06]. Copyright [PBB02]. CORBA [DPPR03, OOS01a, OOS01b, Vit01]. Cornucopia [Che05a]. Corporate} [Bar03]. Cost [Got08l]. Could [Cos04d, Got08h]. Coupled [LMH05]. Course [BV07, Kul07, Pet08b, Pet08c, Yeh06, Yeh07]. Covers [vS01]. Coyle [Puz02a]. Cradle [Got06f]. Craig [Vol02]. Crash [IM01]. Creating [DSD07]. Crisis [Got08a]. Critical [Sto06a]. Critics [Got04a]. Crossings [Got08h]. Cryptography [PBB02]. Curbing [Sto03b]. Cure [Cos03b].

Curious [Puz02b]. Current [Xue02a]. Curricula [Kor08, Pap06]. Curriculum [Bra07, BM07, MSC^{+}02, SMN^{+}02]. Cyber [Dwi02a, ET04a, HAMC^{+}04]. Cyber-Terrorism [ET04a].

D [ET03]. Dagum [Stp02]. Dale [C.03]. DAML [MFHS02, Shi01]. Dan [Bar03]. Dark [Vin02]. DARPA [Pow01]. Data [Abr03b, Ano03, Att07, CSZ03, DFCP03, Gam06, GGC03, Got06g, Got06i, Got07e, LHWX03, Lin03a, McG03c, Meh02, OT07a, Rod02, SRJ06, YPSW04, KLZ04, RH04b]. Data-Centric [GGC03]. Data-Driven [Got06g]. Data-Gathering [Rod02]. Data-intensive [YPSW04, KLZ04]. Data-Storing [McG03c]. Database [BB02, Kim03, Var06]. Datamart [Ano03]. Dave [Stp02]. David [Abr03a, Iva03, Kim03]. DAVIM [HMJV08]. Days [App02c]. DDoS [Got07g]. Debate [Dav03, Got04b]. Decision [Abr03b]. Degrees [Lew02b]. Delay [Got06a]. Delay-Tolerant [Got06a]. Deliver [Sta02b]. Delivering [Puz03a]. Delivery [BPBT06]. Demand [Alf06a, Alf04a]. Demands [Sto02e]. Demandware [Alf04a]. Demise [Got06j]. Departure [OT06b]. Dependable [GBK03, MSCV05, NABG03]. Dependency [DMJ^{+}02, Lin03a]. Deployment [JRC^{+}06, Meh02]. derful [Got06h]. Design [BCA^{+}01, Cla03, DFCP03, Pha01, Sem02c, Veg07, Wil06, Mir02]. Designing [AvSQP05, DMJ^{+}02, OOS01a, SRJ06]. Desk [RCK01]. Desktop [Hed00a, Hed01a]. Desktops [SYS07]. Develop [Cha03]. Developing [Abr03a, Cha03, SER^{+}04, She02b]. Development [Abr03a, Ano02b, Ano03, Ber05b, CFJ^{+}02, Got02c, KRG08, Mor03, PJMW08, Pet08a, Sto06b, Zen02]. Deviant [HWCG06]. Device [AR07]. Devices [AHC^{+}07, Bar03, BCS02, STHK02]. Diabetes [DMJ^{+}02]. Dial [Vil03]. Dieter [Sem03b]. Different [Sto06b]. Digital [Cos04b, Got07b, Shn02]. DiMA [DMJ^{+}02]. Directions [KGR02]. Discovering [DB04]. Discovery [DFCP03, FCBG07]. Discussion [Mil02, Nec06, Sto06a]. disk [HGG^{+}04b]. disks [HGG^{+}04a]. Disorder [PSB^{+}03]. Displays [LHWX03]. Distance [Hed01b]. distant [Cos04c]. Distributed [AKS07, Ano00, Bak01, BR02, BV07, BR07, Cai03, Che05b, DvdHD^{+}08, FvdMAR08, FHR^{+}08, FJH^{+}08, FBG^{+}08, FJ07, FH06, Gam06, Iva03, Kim02, Kul07, Lat05b, LCF^{+}08, MSC^{+}02, Mil00b, OSE06, Pap06, Par04, PMF07, Rao08, Sas02a, Sch01a, Sha06, SS06, SGP04, SRJ06, SKA^{+}02, Wil04, Woo02, Xue02a, vS01, AvSQP05, ATS04, vS04, Gil00, Xue02a]. Distributed-Computing [BV07]. Distributed-System [FH06]. DNA [McG03c]. Do [Ugu02, Hil03]. Doctoral [CM06]. Dolev [Jak02]. Done [Pie06]. Donley [Pér03a]. Don’t [App02a]. Door [OT07c]. Dornfes [Veg03]. Douglas [Dwi02b, Puz02b, She02a]. Dournaee [Sem02a]. Down [Got06m]. Draws [McL03b, McL03c]. Driven [Got06g, KRG08, PJMW08, SSD^{+}04]. Drivers [Sto03b]. Dror [Stp02]. Drug [DMJ^{+}02]. DS [Bac02, Sch01c, ST07]. Duane [Zen02]. Dubuisson [Ros03]. During [Dav03, GMG07]. dynamic [RH04a].

E-Business [Sem03b, Sem03b]. e-commerce [Alf04a, Che02, Che02, Gal05]. e-Learn [McG03b]. E-learning [Gan08]. E-Textile [DMJ^{+}02]. E-Voting [Got06f, Alf04b]. Early [App02c]. Earthquake [Gee06]. Easing [Got03b]. ebXML [PN03]. Economics [Cla03]. ed [Feu02, Puz02b]. ed.} [Dwi02b]. Editable [Ber05a]. edited [Sta02a]. Editing [Qin01]. Edition [ET03, Sem02b, She02a, Zen02]. editor [YPSW04, Kor07a, Kor07b, Kor07c, Meh02, Met03, Pie06, Reg06, San02, Sat03, SK06]. Editors [CP03, CSZ03, DF03, Got06b, HH03, LV03, SS06]. eds [Sem03b, Sti03]. Education [FvdMAR08, FHR^{+}08, FJH^{+}08, FBG^{+}08, Kum03, Lie06, LCF^{+}08, MSC^{+}02, Mur06, Pap06, Sto02j]. Effective [RK06]. Elder [Sta02b]. Electronic [Cou00, WR00a, WR00b]. Embedded [DSD07, Mir02, Reg06, SMN^{+}02]. Embedded-Systems [SMN^{+}02]. Emergence [ET04a]. Emerging [Hed01b, KLAR08, PMF07]. empowering [Dav04b]. Enable [Yan05]. Enabled [Puz07c, WSKP01]. End [DFK08, Got04b]. End-to-End [DFK08, Got04b]. Enemy [Cha04c]. enforcement [Dav04b]. Engineering [CL02, Fox03, Got08e, Wea01a, Wil04, Vil03]. Enough [DB04, San03]. Ensure [Got08b]. Enter [Puž07b]. Enterprise [Bar03, C.03, Cha03, Got06g, Iva03, Var06, Mil07c]. Enterprise-Secure [Cha03]. Environment [Epp03, Toh04]. Environments [AKS07, FCBG07, Gam06, Hed01a, PSB^{+}03, Pet05, Puz04, dRE06]. Era [Got02c]. Eric [Mor03]. Essential [OT06b, Zal04]. Essentials [Lin03c]. Estimating [KLZ04]. Etherware [BGSK04]. Event [Iva03, PMF07, SGP04]. Event-Based [PMF07]. Eventizing [FPA05]. Events [Iva03]. EventWeb [Jai07]. Everyone [Man07]. Everything [Puz02b, Ugu02]. Evolution [Kim02]. Evolving [dRE06]. Excellent [Cha05, Kar04a, OT07b, Sas02a]. Executable [GMG07]. expected [Cos04a]. Experience [Abo03, FJH^{+}08]. Experiences [Pet08c]. Expert [Nec06]. Exposed [Cha03]. Extends [Got06l]. Extensible [Ros06]. Extensive [GF08]. Extreme [BAB01].

F [Vol02]. F. [C.03]. faces [Got04b, Cha05]. Facilitating [Ano00]. Fair [Sas02b].

Fairness [Got08b]. Falls [Bak01]. Fast [Got07d]. Faster [Got07c, Pow01]. Father [McG03d]. Fault [GBK03, IM01, Kim02, Qin01, Toh04]. Fault-Tolerant [IM01, Kim02]. FBI [McL03c]. Feedback [PSB^{+}03]. Feeds [Sem07]. Feet [Sta03a]. Fi [Alf06b, Cos04d, Got03c, Got08c, Sto02g]. Fields [OT06c, Zen02]. fight [Cos04b]. File [HWCG06]. File-Sharing [HWCG06]. Filtering [LSY03]. Financial [Got08a]. Find [DMJ^{+}02]. Firewalls [ET03]. Flexible [Yeh06, Yeh07]. Floating [JG08]. focuses [Alf04a]. Fogg [Hil03]. Folk [Jai02]. Food [Sta03b]. Form [Got07b]. Formal [KP06]. Formal-Methods [KP06]. FORMI [KDH^{+}06]. Foster [Bak04]. Foundation [Shi01]. Foundations [Lin03c, Gor04b]. Fragmented [KDH^{+}06]. Fragments [Yeh07]. Framework [LL03, RCS08, VT01]. Frameworks [Ber05b]. Frank [Puz02a]. Fraud [Wea02]. Fraudulent [Dav04b]. Fred [Mil00b]. Free [HCW05]. Friendly [OT07d]. Full [Bak01, Mor03, Puz06b]. Functionality [Got06c]. Fundamentals [BR02, Cha04d, Dwi02b, Gil01]. Furthers [Gee06]. Future [CFJ^{+}02, KGR02, Met03, Mil05, Vol02, Yan06, Cor04, Mil04].

G [App02d, BAG01, Got03a]. Gadgets [Nac03]. Gaining [GMG07]. Games [ST07]. Gaming [Sto02b]. Ganek [Mil07a]. Garg [Kar02, Vil02]. Gasós [Sem03b]. Gathering [Rod02]. Gen [Got04c, Got06l, Got07a]. Generation [Ahm07, Hed01b, Puž07b, KP05]. Geographical [Cur07]. Gerard [Sas02a]. Get [Ahm07, Got06d, McG03a]. Gets [App03, Got06e]. Getting [App02b, Cha04c, Got08k, Sto02c]. Glass [Woo02]. Global [Got06d, Meh02]. Globus [Pet08a]. Gnutella [Got06b, HCW05]. Go [Bar03, Cos03a]. Goals [PHM03]. Goes [Bar03, Sta03a]. Going [Fra03, Wea00].

Goldberg [Sti03]. Goleniewski [Lin03c]. Good [Bin04, Pér03b, San03, SOB02, Zen02]. Google [Veg03, Veg03]. GovNet [Got02b]. GPS [Cos04d]. Graduate [BV07, Kul07]. Graham [Woo02]. Grande [Fox03]. Graphical [ACR01]. Graphics [BCPK03, FJ07, TCB03]. Gray [Mil04]. Great [She02a]. Grid [Col04, AB07, App02b, Bak04, BAG01, BV07, CKB06, DF03, FJ07, Fox02, Gee06, Gen02, Got06e, JRC^{+}06, KLAR08, Lat05a, Lat06a, McL02, MM07, Mur01, Pet05, Pet08b, Pet08c, Sed06, Sto06a, Sto02b, WF06a, WF06b]. Grid-Based [Gee06]. Grids [SRJ06, DPPR03, LMH05, Puz07a, Puž07b, TT03, Wei00, dCKC06]. Grips [Sem07]. Ground [Sed06]. Groups [Cos03d]. Groupware [IM01]. Growing [Got08g]. Grows [DF03]. Guest [YPSW04, CP03, CSZ03, DF03, HH03, Kor07a, Kor07b, Kor07c, LV03, Meh02, Met03, San02]. Guide [Abr06, BBdGM02, Dwi01, Dwi02a, GF08, Mye01, Pet08a, Puz02a, SOB02, Wea00].

H [Puz03c, Ugu02]. H. [Vol02]. Hacker [Bra07, BM07, ET03]. Hackers [Bra07]. Hacking [Cha03]. Hacks [Veg03]. Hand [RCK01]. Handbook [Ano03, Sas02b]. Hands [Fei07, Mye01]. Hands-On [Fei07, Mye01]. Hanging [Fer01]. Hanna [Sem03c]. Hansmann [Sas02b]. Hardening [Man07]. Harder [Got08k]. Harkey [Bar03]. Harper [Abr03a]. Harvesting [Veg03]. Hatcher [Mor03]. Health [Got06d, PSB^{+}03]. Heel [Sat03]. Helps [DMJ^{+}02, Stp02, Cos04b]. Henry [Puz03a]. Here [Got07l]. Heterogeneity [FCBG07]. Heterogeneous [BFHX05, Ros03, dRE06]. Hides [Zal03]. High [Kar02, Mur06, Mur07, OT07c]. High-Performance [Mur06, Mur07, OT07c]. Hit [Got07d]. Hoc [FCBG07, Got02e, SGP04]. hold [Alf04b].

Home [Nec07, Puz04]. hope [Alf04b]. Hosted [Got06i]. Hotels [Cos03a]. House [Got03c]. HTML [Sem02c, Ugu02]. HTTP [Zal03]. Huge [Got07k]. Hugh [Kim03]. Human [Abo03, Abr06, Shn02]. Human-Computer [Abr06]. Humans [Gan06]. Hundred [Sta03a]. Hypermedia [BBdGM02].

Ian [Bak04]. Idea [Cos03c]. Identification [SK06]. Identity [Dav04b, Epp03]. Idle [Cos03b]. IEEE [Abo03, App02c, BCPK03, Bir03, CP03, Cla03, DMJ^{+}02, DF03, Epp03, FH03, Fra03, HH03, Huh03, LV03, LL03, MFHS02, Met03, Nac03, PSB^{+}03, PKP03, PN03, PMP03a, PMP03b, PMP03c, PMP03d, PMP03e, PHM03, Rod02, San02, San03, Sat03, Shn02, Sta03b, Sta03a, TT03, TCB03, Vin03c, Vin03b, Vin03a]. II [WF06b]. Ilya [Kar02]. Imaging [Shn02]. Impact [Nac03]. Implementation [BCA^{+}01, Abr03a, Vil02]. Implementations [Vil02]. Implementing [Sat02b]. Implicit [Kar03]. Importance [Puz07a]. Inadequate [Kow06]. Incorporating [Toh04]. Industrial [Veg03, Sta02a]. Industrial-Strength [Veg03]. Industry [LSY03, PKP03, Sta02a]. influence [Kar04c]. Information [Abr03a, C.03, Got06d, Law04, Var06]. Infrastructure [FBG^{+}08, Man07, TCB03]. Infrastructures [HAMC^{+}04, PHM03]. Initiation [Puz03a]. Innovation [Got08b]. Innovative [BR07]. Insight [GMG07]. Integrated [DMJ^{+}02, GK02, Kim03, Xue02b]. Integrating [KDH^{+}06, Pap06, RSL06]. Integration [Rim02, Vin02, Vin03c, Vin03b, Vin03a, HGG^{+}04a, Pér03a]. Intellectual [LCF^{+}08]. Intelligence [Cha04b]. Intelligent [Cos03c, Gan08, Kar04c, MFHS02, PSB^{+}03, Shi01, Vol02]. Intelligent-Agents-Based [PSB^{+}03]. intensive [KLZ04, YPSW04]. Interaction [Abr06, STHK02, AvSQP05]. Interactive [MSC^{+}02, Sha03, Hed00b]. Interceptors [RSL06]. Interest [SKA^{+}02]. Interleavings [BAB01]. International [CM06, FHR^{+}08]. Internet [Dwi02b, ET03, Law04, Puz02b, Puz03a, Puz03b, Alf05a, BCS02, Bir03, CP03, Cla03, DF03, ET03, Fra03, Got04b, Got08k, HH03, Huh03, JG08, LV03, Meh02, Met03, Mil00a, OT07b, PKP03, PN03, PHM03, Pow01, Puz02b, Puz03b, San02, San03, STHK02, Sch01c, Sem02b, Sti03, Sto03c, TT03, Vin02, Vin03c, Vin03b, Vin03a, Zal03]. Internet-Based [Sch01c]. Internet-HTTP [Zal03]. Internets [Dwi02b]. Internetworking [She02a]. Interview [Ano01, Mil00b, Mil07a, Mil07b]. Introduces [Gil01]. Introduction [Abo03, Bac00, CP03, CSZ03, C^{+}02, DF03, Fox01, Gor04a, HH03, Iva03, Kor07a, Kor07b, Kor07c, Kow06, LV03, Meh02, Met03, OT07d, San02, Sas02a, Sem02b, She02b, Sti03, Woo02, Zen02, Cor04, YPSW04, Sas02a]. Intrusion [ET06]. Invasion [Sto03c]. Investigating [Sto04]. Invocation [Vin03a]. IP [Puz03c, She02a, Fei07, Mye01, Puz03a, Puž06a]. IPTV [Alf05a]. IPv6 [Met03, Puz07c]. Ire [Cos03c]. IS-95 [Vil02]. ISP [Got08b]. Issue [Got08k]. Issues [Kim02, vS01, HGG^{+}04b]. Item [LSY03]. Item-to-Item [LSY03].

J [Hil03, Vil03]. J2EE

[Cha03, Sem03a, AKS07, Sem03a]. JADE [GF08]. Jamalipour [Puz03b]. James [Ugu02]. Jamsa [Sem02c]. Java [Ano02b, Bar03, Cha03, Che02, Mor03, Ano02b, ACR01, BM01, Ber05b, Cha03, Che02, Feu02, Fox03, Gil00, Gil01, Got06j, KDH^{+}06, Par04, Yan05]. JavaServer [Zen02, Cha05, Zen02]. Jeff [Stp02]. Jim [Mil04, Sem03a]. Johnston [Puz03a]. Jorge [Sem03b]. JPL [Sch01b]. JSP [Sem03c, Sem03c]. Jumpstart [Abr03a].

Jury [Wea01b]. Just [Vin03b, HMG05, Puz02a].

Ka-Ching [Got07h]. Ken [Lin03a, Sti03]. Kennedy [Lin03a]. Keogh [Sem03a]. Kevin [Puz03c]. Key [Cos03b, Got07d, PHM03]. King [Sem02c]. Kit} [Pér03b]. Klaus [Sem03b]. Klaus-Dieter [Sem03b]. Knots [Gor07]. Know [Cha04c, Puz02b]. Knowledge [Gor04b]. Knowledgeable [She01]. Knows [App02c]. Kohr [Stp02]. Kolb [Zen02]. Koniges [Sta02a]. Konrad [Sem02c]. Kris [Sem02c]. Kutnick [C.03].

L [Cai03]. Laboratory [Buy01]. Lags [Got06f]. Lamport [Mil02]. Landscape [Got06l]. Lane [Kim03]. Language [Kar03, MFHS02]. Large [Gam06, SRJ06, RH04b]. large-scale [RH04b]. Large-Volume [Gam06]. Larkin [Bar03]. Last [Sta03a]. Latecomer [IM01]. Latency [FH06]. Laurence [Xue02b]. law [Dav04b]. Layman [Cha03]. LDAP [Pér03a, Cor04, Pér03a]. League [Bir03]. Learn [Ano02b, Bra07, McG03b, Pha01, Veg07]. Learning [AS07, BTSV^{+}07, Cha03, Cha05, Hed01b, Puz03c, Ugu02, Gan08]. Legacy [Got06b]. Leonardo [Stp02]. Leslie [Mil02]. Lessons [Cha04b, Mil04]. Lets [Cos03e]. Level [Ram01]. Leveraging [SS06]. Liberty [Far02, Got06f]. Libraries [Got07b]. Library [Sem03a, McL02]. Life [Alf06b, Got06h]. Lillian [Lin03c]. Limit [Sto02i]. limiting [Dav04b]. Line [McL03b, SK06, Alf04a]. Linux [OT07e]. Linux-Based [OT07e]. Liskov [Mil05]. Little [Gor04a, Got04b, Got06k]. Liu [Cai03, Puz03c]. Live [McG03b]. Load [OOS01a, OOS01b]. Local [NABG03]. Locality [Att07]. Location [PMP03a, PMP03b, PMP03c, PMP03d, PMP03e]. Location-Aware [PMP03a, PMP03b, PMP03c, PMP03d, PMP03e]. Locks [ACR01]. Logical [Gor04b]. Long [Got07j]. Long-Term [Got07j]. Look [Col04, Sti03]. Looking [Got06b]. Loosely [LMH05]. Lothar [Sas02b]. Loughran [Mor03]. loved [Cos04c]. Low [Got08l]. Low-Cost [Got08l]. Luckham [Iva03].

M [ET03]. M. [Ano01, Cai03]. Mail [Cos03c]. Mainstream [Got03b]. Maintenance [NABG03]. Majoring [Sto02d]. Making [Got07i, Wea01a]. MAN [Got04c]. Management [AKS07, CSZ03, DFCP03, DMJ^{+}02, Epp03, GGC03, Got07d, Got08a, Got08b, LHWX03, Mil07a, Mil07b, OSE06, Pér03a, Puž06a, Xue01, dRE06, DLBC04, HMG05, RH04b]. Manager [Puz02a]. Managers [Puz02a]. Managing [PBB02]. Mangano [Bon03]. Manual [Kar02]. Manufacturing [Bad07]. Mapping [Vin03b]. Mappings [GMG07]. Mark [Mil07b, Zen02]. Market [Got06k]. Marking [Got02b]. Mart [Ano03]. Martin [Sas02b]. Mashups [Cur07]. Master [BR07]. Mastering [Che05b]. Matter [Vin02]. Maydan [Stp02]. McDonald [Stp02]. Mean [Got07h]. Media [Nac03, HGG^{+}04b, HGG^{+}04a, Kar04c]. media-based [Kar04c]. Meeting [Shn02]. Meets [Got05, Got06c, Kil02]. Melbourne [Tan07]. memory [HGG^{+}04a]. Menon [Stp02]. Merk [Sas02b]. Message [Fox02]. Metaheuristics [Cap08]. Metamodel [PLB07]. Metamodel-Based [PLB07]. Methods [HL07, KP06]. Mexico [Lie07]. Microsoft [Mil04, Opp03, Pér03b]. Middle [Got04b]. Middleware [AR07, Bac04, Ban03a, Ban03b, BCCK01a, BCCK01b, CM06, DvdHD^{+}08, FCBG07, FPA05, HM06, HAMC^{+}04, HMJV08, JRC^{+}06, LV03, Liu07, OOS01b, Paw06, Rim02, RCK01, Ros06, RCS08, SER^{+}04, SWH06, VT01, Vin02, WSKP01, dRE06, DLBC04, KP05]. Middleware-Based [OOS01b]. Might [Cos03b, Got07h]. Migrating [Nac03]. Mike [Bar03]. Miller [Xue02b]. Mining [Abr03b, Wea02, ATS04]. Mirrors [Got07c]. Mobile [Abr06, AHC^{+}07, Ano01, AS07, AR07, Bar03, BTSV^{+}07, CP03, CSZ03, Cos02, DFCP03, FCBG07, GGC03, Got02d, Got07f, HL07, KGR02, Kul07, LHWX03, Mjø02, Nac03, Pra01, Puz03b, Puž06a, SER^{+}04, Sat02a, SWH06, SGP04, Mjø02]. Mobility [Puz06b]. MobSim [LRA01]. MoCA [SER^{+}04]. Model [KRG08, MSCV05, PSB^{+}03, PJMW08, Ano03]. Model-Driven [KRG08, PJMW08]. Modeling [Gee06, Liu06, SKA^{+}02, VT01]. Models [GMG07]. Modern [Lin03a, Pha01]. Modular [VPW^{+}06]. Money [Got07i]. Monitoring [Col04]. Motivating [Jak02]. Moving [DFCP03, Met03, Ano02a]. MPI [Vre04]. Multi [Ahm07, GF08]. Multi-Agent [GF08]. Multi-View [Ahm07]. multiagent [Gue04]. Multidisciplinary [SMN^{+}02]. Multifractal [Liu06]. Multimedia [Lin03b, Puz03a, Nac03, Shn02]. Multiple [HGG^{+}04a]. Multiprocessor [SBR^{+}07]. Multipurpose [Nac03]. multithreaded [Omm04]. Multiuser [SS06]. Multivendor [RK06]. Museum [Hed00b]. Music [Hed00b]. my [Alf05a]. MySQL} [Kim03]. Mystery [Sch01a].

Nanorover [Sch01b]. Napster [Got06b]. Narrow [Xue01]. NASA [Sch01b]. NATs [Tou02]. Navigability [Rod02]. Nee [Lin03b]. Need [Got04a, Puz02b]. Needs [Shn02]. Net [Got07f, Got07i, McL03a]. Network [EA07, Gil01, Got06a, Got06c, Man05, Man07, Mid02, Pér03b, Puz07a, Puz07c, C.03, Pér03b]. Network-Centric [Mid02]. Networked [BGSK04, BJ00, DFCP03]. Networking [Bra07, Dwi02b, Got08g, Mye01, Nec07, Puz02b, Puz04, She02a]. Networks [BPBT06, CKB06, Dav04a, Dwi02b, GGC03, Gon02, Got06d, Got08d, HM06, HMJV08, HWCG06, LRA01, McL06, NABG03, Puz06b, SGP04, Tan07, Yan06, BFHX05, Sto04, Lin02]. Newest [ST07]. News [App02c, App02d, App03, Got02d, Got03a, Got03b, Got03c, Got06f, Got06g, Got06k, Got06i, Got06j, Got06h, Got07e, Got08i, McG03c, McG03d, McL03a, McL03b, McL03c, Sto02g, Sto03c]. nex [KP05]. nex-generation [KP05]. Next [Ahm07, Got04c, Got06l, Got07a, McL04]. Next-Gen [Got04c, Got06l, Got07a]. Next-Generation [Ahm07]. Nicklous [Sas02b]. Nikhil [Kar03]. Nimrod [BAG01]. Nimrod-G [BAG01]. No [Fer01]. Norm [HWCG06]. North [WF06a, WF06b]. Notification [SGP04]. Novices [Stp02]. Numerical [Dav04a, Pet05]. Nutshell [Che02]. NUTSS [Fra03].

Objects

[ACR01, DPPR03, KDH^{+}06, Vit01].

Obsessive [PSB^{+}03].

Obsessive-Compulsive [PSB^{+}03]. Ocean [JG08]. ODAM [ATS04]. Oecumenism [KP06]. OFDM [Lin03b, Lin03b]. Offers [Got07k]. Oh [Got03a, Got08f]. OIL [MFHS02]. ok [Got04b]. Old [Got07k]. Olivier [Ros03]. OMA [Sto02h]. On-Demand [Alf06a, Alf04a]. One [Dwi02b, Sto02a]. One-Stop [Dwi02b]. ones [Cos04c]. Online [Dav04b, FJH^{+}08, SSD^{+}04, Sti03, ST07, Bac02]. Ontology [MFHS02]. Open [Alf05b, Cla03, Got05, Sto02e, Veg06, BCA^{+}01]. Open-Sourcing [Sto02e]. Opening [Got07f]. OpenMP [Stp02, Vre04]. Operating [Xue02a]. Opponent [Cos03c]. Opportunistic [dCKC06]. Opportunities [Got04c, Mjø02]. Optimization [Lin03a]. Optimized [RH04b, ATS04]. Optimizing [Kar02, Lin03a]. ORB [BCA^{+}01]. Oriented [Bon06, BAG01, Sto06b]. Outlook [Opp03]. Outposts [Sch01b]. Overcome [FCBG07].

Overcrowded [Got02a]. Overview [Yeh06]. Owner [Kar02]. Ownership [Dwi02a]. Oxygen [Hed00a].

P [Kar02, Puz02a]. P2P [App03, HWCG06, Kil02, TT03, Wea01b]. Pages [Sem02c, Zen02, Zen02]. Papers [Ban03a, Ban03b]. Parallel [BJ00, Cap08, Cha04a, DPPR03, Fox02, Kar03, Kar04a, Lat05b, LRA01, Mar08a, Pap06, Pet05, Sta02a, Xue02b, BJ00, Stp02, Vre04]. Part [DFCP03, FvdMAR08, FHR^{+}08, FJH^{+}08, FBG^{+}08, GGC03, LCF^{+}08, Pet08b, Pet08c, HGG^{+}04b, WF06a, WF06b]. Partners [Bac02]. Passing [Fox02]. Past [Got08c, Yan06]. Patents} [Dwi02a]. Patients [Got08j]. Patterns [RCS08, Sem06]. Paul [Ano03]. Payment [Wil06]. PC [Kil02]. PCS [Vil02, Cos03b, Got08l]. Peace [PO01]. Pedestrian [Gav01]. Peek [Vol02]. Peer [Bir03, Cla03, Fra03, Gon02, Got06b, HWCG06, SRJ06, TT03, BFHX05]. Peer-to-Peer [Got06b, HWCG06, SRJ06, BFHX05]. Peering [BPBT06, Kil02]. Peers [Wea01b]. Pence [Ugu02]. Pentagon [Sto02i]. People [PSB^{+}03]. PeopleVision [PSB^{+}03]. Performance [FH06, Kar02, Mur06, Mur07, OT07c, Vin03c]. Perils [Got06i]. Perl [Che05a]. Personal [App02c, LRA01]. Perspective [Bon06, Rim02]. Perspectives [Sto06b]. Persuasive [Hil03]. Pervasive [Abo03, Ano01, App02c, DMJ^{+}02, Epp03, FH03, Kum03, PSB^{+}03, PMP03a, PMP03b, PMP03c, PMP03d, PMP03e, Sat03, Sta02b, Sta03b, Sta03a, TCB03, Yan05, Sas02b]. Pesky [Tou02]. Peter [Vol02]. pH} [Kar03, Kar03]. PhD [Pet08b, Pet08c]. Phil [Sem03c]. Philosophical [Gor04b]. Phone [ZN06, Cos04a]. Phones [Cos02, Cos03e, HL07]. PHP [Kim03]. Pinstripes [App03]. Pipelines [FJ07]. Piracy [Cos03d]. Placement [SSD^{+}04].

Platform [FPA05, MSPB02, SBR^{+}07, Alf04a]. Platform-Based [SBR^{+}07]. plus [OT07a]. PoDSy [GBK03]. Point [OT06b, Pér03b, SOB02]. Policies [RSL06, Sto02i, VPW^{+}06]. Politics [Got07g]. Portability [Cos03e]. Portable [BCS02, RSL06]. Pose [TCB03]. Pose-Aware [TCB03]. Posters [Sha03]. postmark [Cos04b]. Potential [Bak01, Got07k]. Potts [Mil07b]. Power [Col04, Got08j, Hil03, Iva03, Sem03c, Sem06, Veg03]. Powerlines [McL03a]. Practical [BR02, ET06, Mor03, Pha01, She02b]. Practice [Xue02a]. Practices [Pet08b]. Prasad [Lin03b]. predicted [Alf05a]. Prediction [Gee06]. Premature [Got06j]. Present [Yan06]. Presentation [HL07]. Presumption [Vin03c]. Prevention [ET06]. Primer [OT06c]. Principle [Att07, Got04b]. Principles [Cai03, Mir02, She02a]. Privacy [Dwi02a, PSB^{+}03, Sat03, Dav04b]. Probabilistic [Ram01]. Problem [Sto02e, Vin03b, HMG05]. Process [Got08c]. Processing [Iva03, Paw06, RCS08]. Production [GK02]. Products [FH03]. Professional [Opp03, Che02]. Professionals [She02a]. Profile [SSD^{+}04]. Profile-Driven [SSD^{+}04]. Profit [Sch01c]. Profits [Cha04b]. Program [Alf06b, BR07, Lie07, Sem02b]. Programmers [Kar04b]. Programming [Ano02b, AB07, Bar03, BM01, Bon06, BJ00, Dav04a, DPPR03, Gil00, Kar03, Kul07, Lat06b, Mar08a, MSCV05, Mil05, Mir02, Par04, Pér03a, Stp02, Veg02a, Vre04, Ano02b, Gil01, Veg02a]. Progress [Bac04, Ban03a, Ban03b, CM06, DMJ^{+}02, Hed01a, Liu07, MCB^{+}06, PSB^{+}03]. Project [Abr03a, Sch01b, Sto06b, Buy01, Hed00a, PO01]. Projects [BCPK03, Sch01a, TCB03]. Property [Dwi02a, LCF^{+}08]. Protect [C.03]. Protection [Gav01, Man05, PSB^{+}03]. Protein [Sch01a]. Protocol [Puz03a, SWH06, Zal03, Alf05a]. Protocol-Based [SWH06]. Protocols [Puz03b, She02a, Puz04]. Prototype [PLB07]. Prototyping [Kor07a, Kor07b, Kor07c, SBR^{+}07]. Provide [PSB^{+}03]. Provisioning [BCS02]. Proxies [RdCVM07]. Public [PHM03, SKA^{+}02]. Public-Interest [SKA^{+}02]. Publishing [Cou00]. Pump [Sha03]. Puts [Sta03b]. Putting [Bad07, Sch01c].

QoS [WSKP01]. QoS-Enabled [WSKP01].

R [ET03]. Racing [Got08c]. Rael [Veg03]. Raising [Cos03c]. Rajat [Kar02]. Ramesh [Stp02]. Ramjee [Lin03b]. Randy [Cha03, Lin03a]. Rapid [Kor07a, Kor07b, Kor07c]. RDF [Gor04a]. Reaches [Sta02a]. Reading [HL07]. Ready [Ahm07, LL03]. Real [BCPK03, LL03, MSCV05, Qin01, Reg06, SMN^{+}02]. Real-Time [LL03, MSCV05, Qin01, Reg06, SMN^{+}02]. Reality [DMJ^{+}02]. Reasoning [VT01]. Recess [Dav03]. Recommendations [LSY03]. Recommended [OT07e]. Recovery [IM01]. Redraw [Got08l]. Reengineering [AKS07]. Reference [Bin04, Dwi02b, Sem03a, She01, She02a, Sem03a, Sem03c]. Reflective [BCCK01a, BCCK01b, RCK01, VT01, WSKP01]. Regs [Dwi02a]. Regulation [Got08b]. Regulators [McL03c]. Reinventing [Gam07]. Relationship [Epp03]. Relevant [vS01]. Reliability [DFK08, Reg06]. Remote [Rod02]. Repelling [ET03]. Replicated [RdCVM07]. Replication [SRJ06]. Report [GBK03, Hed01a, LSY03, PKP03]. Representation [Gor04b]. reputation [RH04a]. Requirements [HGG^{+}04b]. Research [Ano00, Got02c, Got07j, Hed01a, KGR02, SKA^{+}02]. Researchers [Got07a]. Resilient [GGC03]. Resource [BAG01, Kar04a, Kar04b, Lat05b, OT07e, DLBC04]. Resources [DFCP03, Stp04, Mar05]. Respect [App03]. REST [Got04a]. Return [DFK08]. Review [Abr03a, Abr03b, Ano03, Bak01, Bar03, Ber04, Bon03, BJ00, C.03, Cai03, Cha03, Cha04a, Dwi01, Dwi02a, Dwi02b, EA07, ET03, ET04b, ET06, Gan07, Gil00, Gor04b, Hil03, Iva03, Kar02, Kar03, Kim03, Kow06, Lat06a, Law04, Lin02, Lin03a, Lin03b, Lin03c, Mar08b, Mor03, Mye01, OT06a, OT07a, Pér03a, Pér03b, Puz02a, Puz03a, Puz03b, Puz03c, Puz04, Ros03, Sem02a, Sem03b, Sem03a, Sem03c, Sti03, Veg03, Vil03, Vol02, Vre04, Woo02, Zal03, Ano02b, Ano02a, Che02, Feu02, Jak02, Mir02, Mjø02, Puz02b, Sas02a, Sas02b, Sem02b, Sem02c, She02a, She02b, SOB02, Sta02a, Stp02, Ugu02, Veg02a, Vil02, Xue02a, Xue02b, Zen02]. Reviews [Abr06, Bon06, Gan06, Kac07, Lat06b, Liu06, Nec06, OT06b, OT07b, Puž06a, Puz07a, Ros06, Sed06, Sto06b, Wil06, Yan06]. Revisited [HCW05]. ReX [Ano00]. RFID [McG03d, Sta03a]. Richard [Abr03a, Lin03b]. Riding [HCW05]. Rights [LCF^{+}08]. Ring [Got07h]. Ring-Ring [Got07h]. Rise [ZN06]. Rishiyur [Kar03]. Risk [Gee06, Got08a]. Risk-Management [Got08a]. Risks [Cha03]. RMI [KDH^{+}06]. Road [Sto02c]. Rob [Veg02a]. Robots [Sti03]. Rohit [Stp02]. Roland [Sti03]. Role [Got08a]. Ross [Vil03]. Royalties [Got07i]. RSS [Sem07]. Rubin [ET03]. rule [ATS04]. Rules [Dwi02a]. Runtime [GMG07]. Russ [Xue02b].

S [Kar03, Sas02b]. Safeguard [Sto02h]. Safety [App02c]. Sal [Bon03]. Sales [GK02]. Sanctions [Dav04b]. Sans [Lin02]. Satisfy [PHM03]. Satyanarayanan [Ano01]. Say [Got03a, Got04a]. Scalable [DFCP03, Stp04]. scale [RH04b]. Scheduling [Ram01]. Schneider [Mil00b]. Scholarship [Lie06, Lie07]. School [Bar03, FJH^{+}08]. Schools [FHR^{+}08, Mur07]. Science [Fox03, MSC^{+}02, Pap06]. scientific [RH04a]. Script [Got06j]. Scrutiny [McL03c]. Search [Got07e, McL04, BFHX05]. Searches [Lew02b]. Seattle [Hed00b]. Second [ET03]. Secret [NABG03]. Secrets [Ros03]. Secure [C.03, Cha03, Epp03, Feu02, Sto03c]. Securing [C.03, HAMC^{+}04, Sem02a]. Security [Cha04d, EA07, ET03, Far02, GBK03, Got02e, Got06c, Opp03, PSB^{+}03, PHM03, RSL06, San03, SK06, Vil03, Wil06, Sem02a, ET04b, Feu02]. Seeing [Abr03b]. Seidman [Abr03b]. Self [HMG05, Jak02, KRG08, Jak02]. Self-Adaptive [KRG08]. Self-Stabilization [Jak02, Jak02]. Semantic [Law04, OSE06, PSS02, VT01, Got06g, MFHS02, vH04]. Sensible [Man05]. Sensor [CKB06, Gav01, GGC03, HM06, HMJV08, NABG03, Tan07, Yan06]. Sensor-Based [Gav01]. Sentient [Huh03]. Sequential [Xue02b]. Server [Abr03b, Ano02b, Ano02b]. Server-Based [Ano02b]. Server-Side [Ano02b]. Servers [AKS07, Ram01]. Service [BGSK04, BAG01, DFCP03, FCBG07, Got08e, Lie06, Lie07, Mil07a, Mil07b, OOS01a, SRJ06, Sto06b, VPW^{+}06]. Service-Oriented [BAG01, Sto06b]. Services [BCS02, C^{+}02, DB04, ET04b, GK02, Got03b, Got04a, Kow06, LV03, PKP03, PN03, RdCVM07, SSD^{+}04, Woo02, Puz03b]. Session [Puz03a]. sets [RH04b]. Shan [Bar03]. Sharapov [Kar02]. Sharing [Gam06, HWCG06, Var06]. Shirts [App03]. Shlomi [Jak02]. Short [Bak01]. Should [Got03a]. Side [Ano02b]. Siegwart [Sti03]. Sign [Got08d]. Sign-on [Got08d]. Signals [Got04c]. Simmers [Dav03]. Simple [Pér03a, vS04]. Simplifying [Gam06]. Simulation [LRA01, ST07]. Single [Got08d, SS06, HGG^{+}04b]. single-disk [HGG^{+}04b]. Single-User [SS06]. Sinnreich [Puz03a]. SIP} [Puz03a]. Sites [Got08g, Veg07]. Six [Lew02b]. Skies [Col04]. Slouching [Got06g]. Small [Got06e, LHWX03, STHK02]. Smallpox [Cos03b]. Smart [Hed01a, RdCVM07, ZN06]. Smith [Vol02]. Smoke [Got07c]. SOAP [C^{+}02]. Social [Got08g, Got08d, Wil04]. Software [Gan07, Got08a, Got08e, Moo01, Rod02, RCS08, Veg06, Kar04c, LL03]. Software-as-a-Service [Got08e]. Solution [MM07, HMG05]. Solutions [Sem03b]. Some [App03]. Soon [Gor04a, Got07h]. Source [Alf05b, Got05, Veg06]. Sourcing [Sto02e]. Space [Wea01a]. Spam [Cos02, Dav03, McL03b, Sto03a, Cos04b]. Spark [Got08e]. Special [FvdMAR08]. Specific [VPW^{+}06]. Speeding [Moo01]. Spoon [Paw06]. Spotlight [PN03, PMP03a, PMP03b, PMP03c, PMP03e, ZN06, PMP03d]. Spyware [Cos03d]. SQL [Abr03b]. Stabilization [Jak02, Jak02]. Stack [Got06m]. Stallings [Lin02]. Standard [Got04c, Got06k, Sto02f]. Standards [Got07d, Got08c, Sem07, Ber04]. Starting [Pér03b, SOB02]. State [Gan07, OT07c, McL03c]. State-of-the-Art [OT07c]. Statecharts [DSD07]. Stepping [Got07j]. Steve [Mor03]. Steven [ET03]. Still [Got02b, Got02e]. Stober [Sas02b]. Stop [Dwi02b, Gam07]. Storage [GGC03, HGG^{+}04b, HGG^{+}04a, Sto02a, dCKC06, RH04b]. stored [RH04b]. Storing [McG03c]. Strategies [Ano02a, C.03, OOS01b, dCKC06]. Strategy [Abr03a]. Streamline [Got08h]. Strength [Sta02a, Veg03]. Strike [Lew02a]. Structure [Got07e, Pet08c]. Students [AB07, Pet08b, Pet08c]. Study [GF08].

Styles [Vin03a]. Submissions [vS00]. Success [Sas02b]. Suitable [Man07]. Summary [Liu06]. Summer [FHR^{+}08]. superhighway [Mar05]. SuperNet [Pow01]. SuperNets [Bir03]. Supply [Alf04a]. Support [IM01, Qin01, Vit01, HGG^{+}04b, HGG^{+}04a]. Supporting [VPW^{+}06]. Surveillance [DvdHD^{+}08, PSB^{+}03]. survey [DLBC04]. Symposium [CM06]. Synergy [TT03]. System [FH06, Kor07a, Kor07b, Kor07c, Mir02, PLB07, Sat02b, SBR^{+}07, HGG^{+}04b, HGG^{+}04a]. System-on-Chip [SBR^{+}07]. Systems [Abr03a, Bak01, BR02, Cla03, Cou00, DMJ^{+}02, GBK03, GF08, Kul07, KLAR08, LL03, MFHS02, Mid02, Mir02, NABG03, OSE06, Pie06, PMF07, Qin01, RK06, Rao08, Reg06, SMN^{+}02, Sta03a, Veg02b, Vil02, Wil04, Xue02a, vS01, AvSQP05, Gue04, Kar04c, RH04b, vS04, Iva03, Ros03].

T [App03]. T-Shirts [App03]. Table [Sta03b, Cos04c]. Tackle [Sch01a]. Tackles [Alf06a]. Tagging [WR00a, WR00b]. Take [Got03c]. Takes [Mur01]. Taking [Got07b]. Talking [Ber04]. TAO [OOS01a]. Tara [Veg03]. Target [Cos03d, Cos03e]. Taxes [Got07i]. Taylor [Cha03]. TCP [She02a, DFK08, Fei07, Mye01]. TCP/IP [She02a, Fei07, Mye01]. Teaching [AB07, BM07, Fei07, Mar08a, Pet08b, Pet08c, Reg06, Tan07, Veg02b, WF06a, WF06b, Yan07]. Tech [Lie07]. Technical [McG03a]. Technique [Gee06]. Techniques [BJ00, Che05a, Kar02, Xue02a, Sem02c]. Technologies [Bin04, Gor04a, Got06a, MM07, Nec06, Pet08b, Pet08c, Sem02b, Sem03b, Shn02, Wil06, Mjø02]. Technology [BB02, Gan07, Got07k, Got08c, Hil03, Kar04b, Lat06a, McL02, Mil00a, Mur01, San02, Cha03]. Tel [Sas02a]. Telecom [Lin03c]. Telecommunications [Lin03c]. Teleconferencing [BCPK03].

Telemarketers [Cos03e]. Telephony

[Nac03]. Television [Ahm07, Alf05a]. Term [Got07j]. Terminology [Got06c]. Terms [LCF^{+}08]. Terror [Got08k]. Terrorism [ET04a]. tertiary [RH04b]. Testing [BBdGM02, Rod02]. Tests [McL02]. Textbook [Bak01]. Textile [DMJ^{+}02]. Their [DMJ^{+}02]. Theory [Gal05]. Think [Hil03, Mar08a]. Thoben [Sem03b]. Thomas [Sas02b]. Thorough [Sti03]. Those [Pie06, Tou02]. Threads [ACR01]. Throughput [FH06]. Timber [GK02]. Time [Got02b, LL03, MSPB02, MSCV05, Qin01, Reg06, SMN^{+}02, Sto02a, Paw06]. Time-Triggered [MSPB02]. Times [Pow01, KLZ04]. Tips [BM01, Sem02c, Veg03]. Today [Mar08a]. Together [Got06a, Cos04c]. Tolerance [GBK03, Toh04]. Tolerant [Got06a, IM01, Kim02, Qin01]. Tolls [HCW05]. Tomorrow [Sem03b]. Too [Gor04a]. Tool [Cha05, She01]. Toolkit [Pet08a]. Tools [BM01, Ber05b, Veg03]. Topic [ST07]. Tour [BR02, BBdGM02]. Track [Got07d, PHM03, San03]. Traditional [KLAR08]. Traffic [Got08b, Liu06]. Training [FBG^{+}08, Pér03b]. Transformations [PLB07]. Transmission [Puz04]. Treating [DMJ^{+}02]. TRECC [Alf06a]. Trees [Abr03b]. Trends [Vil02]. Trevor [Abr03a]. Triggered [MSPB02]. Troubleshooting [Puž06a]. Trust [RH04a, San02]. Turf [App02d]. Turn [App02a]. Tutorial [OT07b]. Twisted [Gor07].

Ubicomp [PSB^{+}03]. Ubiquitous

[BCS02, CFJ^{+}02, CP03, Epp03, HAMC^{+}04, Sat02a, WR00a, WR00b, Yan07, dRE06]. UDDI [C^{+}02]. Uh [Got08f]. Uh-Oh [Got08f]. Ultra [Got08l]. Undergraduate [Lie07]. Unified [Xue02b]. University [Tan07]. Unleashing [Sem03c]. Unraveling [C^{+}02]. Unstructured [Got07e, BFHX05].

Urban [AHC^{+}07]. Usability [NR02, Rod02]. Usage [CL02]. Usage-Centered [CL02]. Use [BM07, Sto02i]. Useful [Kow06, Lat05b, She02b]. User [BBdGM02, SS06, She01]. Users [DMJ^{+}02, Nec07, SER^{+}04, Veg06]. Using [Ano03, BGSK04, BJ00, Hil03, Kar04b, OOS01a, PLB07, Pet08b, Puz03a, STHK02, Sta02b]. Utilities [Col04]. Uwe [Sas02b].

Validating [DSD07]. Valuable [Kar04b, Stp04]. VB.Net [Ano02a, Ano02a]. Vector [BR02]. Vendor [Gen02]. Venture [Got05]. Verification [PLB07]. Very [STHK02]. via [RSL06]. Video [Ahm07, DvdHD^{+}08]. Vidgen [Abr03a]. View [Ahm07, Xue01, Xue02b]. Views [Shn02]. Vijay [Vil02]. Virginia [McL03b]. Virtual [Cos04c, Buy01]. Virtualization [Got07k]. virus [Cos04b]. Viruses [App02a]. Vision [Gen02, LMH05]. Visions [Shn02]. Visual [PSB^{+}03, Ano02a]. Visualization [ACR01, Law04]. Visualizing [Law04]. Voice [Puz03a]. Voilà [Puz07c]. VoIP [McL03c, OT07d]. Volume [Gam06]. Voting [Got06f, Alf04b]. VR [BCPK03, TCB03]. vs [FH06]. Vulnerabilities [Got06c].

Wal [Ano03]. Wal-Mart [Ano03]. want [Alf05a]. WAP [She02b, Dwi01, She02b]. War [Sto03b]. Warehousing [Ano03]. Wars [App02d, Got08l]. Watches [Vot03]. Watermark [PBB02]. Wave [Lat06b]. Way [DMJ^{+}02, Pha01]. Wayne [Mir02]. WCDMA [App02d]. WDM} [Puz03c, Puz03c]. Wearable [DMJ^{+}02]. Weaving [Sto02g]. Web [Abr03a, C^{+}02, Dwi02a, Kim03, Law04, Puz02a, Sem02b, Sem02c, Woo02, Vol02, Abr03a, BB02, Ber05a, Ber05b, CL02, C^{+}02, ET04b, Fox01, Fox03, Got03b, Got04a, Got06g, Got08f, HH03, Huh03, Kil02, Kim03, Kow06, LV03, LHWX03, MFHS02, McL04, Mil07c, NR02, OSE06, Opp03, PKP03, PN03, PSS02, Pie06, Puz02a, Ram01, RdCVM07, STHK02, Sem02b, Sem02c, Sem06, SYS07, She02b, Shi01, Sto06b, Sto03c, Veg02b, Veg06, Veg07, Woo02, Zen02, vH04]. Web-Based [Veg02b]. Web-Integrated [Kim03]. Webcams [Sti03]. Webtops [SYS07]. Went [Got06k]. Westerman [Ano03]. Wheel [Gam07]. Where [App02c, Kil02, McL06, Vot03, Wea00, Mar05]. Whisper [NABG03]. White [Got03c]. Who [Pie06, Puz02b]. WHYNET [Sto04]. Wi [Alf06b, Cos04d, Got03c, Got08c, Sto02g]. Wi-Fi [Alf06b, Cos04d, Got03c, Got08c, Sto02g]. Wide [Sem02b, Ulm04]. Will [Got03c, Got08b, Got08e]. William [ET03, Lin02]. Williams [Kim03]. Wily [ET03]. Windows [Veg06]. winner [Alf05a]. Winter [FJH^{+}08]. Wired [McL06, Puz04, Rim02, PO01]. Wired-Wireless [Rim02]. Wired/Wireless [Puz04]. Wireless [App02c, Bar03, Bin04, Cos03a, GGC03, Got02e, Got04c, HM06, Kil02, Lin02, Lin03b, McL06, Nec06, Nec07, PKP03, Puz02a, Puz03b, Puz04, Rim02, STHK02, She02b, Sto02d, Sto02h, Sto02i, Sto02j, Tan07, Vol02, Yan06, Sto04, Sto02f]. Wires [Lin02]. Wisdom [FH06, Sha06, SRJ06, Ulm04]. WLAN [Got06h]. WLAN-derful [Got06h]. Wolf [Mir02]. Wood [Abr03a]. Wood-Harper [Abr03a]. Woods [Abr03b]. Work [Bad07]. Works [Bac04, Ban03a, Ban03b, CM06, DMJ^{+}02, Liu07, MCB^{+}06, PSB^{+}03, Puz02b]. Workshop [GBK03]. Workstations [BJ00]. World [BCPK03, Puz03b, Sem02b, Ulm04, Woo02]. Writing [Feu02]. Wrong [Sha06]. WSDL [C^{+}02]. WWW [Mar05].

XML [Law04, Sem02a, Got06m, Sem02a, She01, SOB02]. XML-based [Law04]. XSLT [Bon03, Bon03].

Z [Lat06a].  Zen [HH03].  Zoellick

[Dwi02a].

References

Apon:2007:TCG

[AB07] Amy Apon and Matt Baker.

Teaching Condor Grid computing to beginning programming students. IEEE Distributed Systems Online, 8(4):??, April 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/04/o4002.pdf.

Abowd:2003:IPC

[Abo03] Gregory D. Abowd. IEEE pervasive computing: Introduction: The human experience. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0306/f/b2int.htm.

Abrahams:2003:BRJ

[Abr03a] Alan S. Abrahams. Book review: Jumpstart your Web development project: a review of Developing Web Information Systems: From Strategy to Implementation by Richard Vidgen, David Avison, Bob Wood, and Trevor Wood-Harper. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/d/bks_a.htm.

Abrahams:2003:BRS

[Abr03b] Alan S. Abrahams. Book review: Seeing the woods for the decision trees: a review of Data Mining with SQL Server 2000 by Claude Seidman. IEEE Distributed Systems Online, 4(1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0301/d/bks_c.htm.

Abramson:2006:BRG

[Abr06]  Myriam Abramson.  Book

reviews: a guide to mobile human-computer interaction. IEEE Distributed Systems Online, 7(10):??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/10/ox006.pdf.

Attali:2001:GVJ

[ACR01] Isabelle Attali, Denis Caromel, and Marjorie Russo. Graphical visualization of Java objects, threads, and locks. IEEE Distributed Systems Online, 2(1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0101/features/att0101_
print.htm.

Adamczyk:2007:UCM

[AHC^{+}07] Piotr Adamczyk, Kevin Hamilton, Alan Chamberlain, Steve Benford, Nick Tandavanitj, Amanda Oldroyd, Kate Hartman, Kati London, Sai Sriskandarajah, Eiman Kanjo, Peter Lanshoff, Kaoru Sezaki, Shin’ichi Konomi, Muaz A. Niazi, Hafiz F. Ahmad, Fauzan Mirza, Arshad Ali, George Roussos, Dikaios Papadogkonas, and Mark Levene. Urban computing and mobile devices. IEEE Distributed Systems Online, 8(7):??, July 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/07/
mds2007070002.pdf.

Ahmad:2007:MVV

[Ahm07]  Ishfaq Ahmad.  Multi-view

video: Get ready for nextgeneration television. IEEE Distributed Systems Online, 8 (3):??, March 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/03/
o3006.pdf.

Abdellatif:2007:RJS

[AKS07] Takoua Abdellatif, Jakub Kornas, and Jean-Bernard Stefani. Reengineering J2EE servers for automated management in distributed environments. IEEE Distributed Systems Online, 8 (11):??, November 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/11/
                                        mds2007110001.pdf.

Alfonsi:2004:SDN

[Alf04a] B. Alfonsi. Supply and demandware: New on-demand ecommerce platform focuses on the bottom line. IEEE Distributed Systems Online, 5(12): 2, December 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        iel5/8968/30113/01381276.
                                        pdf?isnumber=30113&prod=JNL&
                                        arnumber=1381276&arSt=+2&
                                        ared=+2&arAuthor=+Alfonsi%
                                        2C+B.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30113&arnumber=1381276&
                                        count=4&index=1.

Alfonsi:2004:VAH

[Alf04b] B. J. Alfonsi. E-voting advocates hold out hope. IEEE Distributed Systems Online, 5(3): ??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/28687/01285879.pdf?isnumber=
                                        28687&prod=JNL&arnumber=1285879&
                                        arSt=&ared=&arAuthor=Alfonsi%
                                        2C+B.J.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=28687&arnumber=1285879&
                                        count=7&index=2.

Alfonsi:2005:WMI

[Alf05a] B. Alfonsi. I want my IPTV: Internet Protocol television predicted a winner. IEEE Distributed Systems Online, 6(2): ??, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/30522/01407761.
pdf?isnumber=30522&prod=JNL&
arnumber=1407761&arSt=&ared=
&arAuthor=Alfonsi%2C+B.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
30522&arnumber=1407761&count=
5&index=2.

Alfonsi:2005:OSC

[Alf05b] Benjamin Alfonsi. Open source in the classroom. IEEE Distributed Systems Online, 6(6): ??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2005/06/o6003.pdf.

Alfonsi:2006:BTT

[Alf06a] Benjamin Alfonsi. In brief:

TRECC tackles on-demand computing. IEEE Distributed Systems Online, 7(3):??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/03/o3007.pdf.

Alfonsi:2006:WFC

[Alf06b] Benjamin Alfonsi. Wi-fi certification program aims to boost battery life. IEEE Distributed Systems Online, 7(1):??, January 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/01/o1004.pdf.

Anonymous:2000:RFD

[Ano00]  Anonymous.  ReX: Facil-

itating distributed research. IEEE Distributed Systems Online, 1(2), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds200/ds2rexprint.
                                        htm.

Anonymous:2001:IMS

[Ano01] Anonymous. Interview: M. Satyanarayanan on mobile and pervasive computing. IEEE Distributed Systems Online, 2 (6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0106/departments/int0106_
                                        print.htm.

Anonymous:2002:VBV

[Ano02a] Anonymous. From Visual Basic to VB.Net: a review of Moving to VB.Net: Strategies, Concept, and Code. IEEE Distributed Systems Online, 3(2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0202/departments/bks_
                                        a.htm.

Anonymous:2002:LAJ

[Ano02b]  Anonymous.  Learn about

Java server-side development and programming: a review of Server-Based Java Programming. IEEE Distributed Systems Online, 3 (2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0202/departments/bks_
b.htm.

Anonymous:2003:BRD

[Ano03] Anonymous. Book review: a datamart development handbook: a review of Data Warehousing: Using the Wal-Mart Model by Paul Westerman. IEEE Distributed Systems Online, 4(4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/0304/d/bks_c.htm.

Applewhite:2002:DTY

[App02a] Ashton Applewhite. Don’t turn your back on computer viruses. IEEE Distributed Systems Online, 3(1), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0201/features/new_print.
htm.

Applewhite:2002:GG

[App02b] Ashton Applewhite. Getting the Grid. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0205/departments/new_
a.htm.

Applewhite:2002:IPC

[App02c] Ashton Applewhite. IEEE pervasive computing: News: What knows where you are? personal safety in the early days of wireless. IEEE Distributed Systems Online, 3(12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0212/d/b4news_print.htm.

Applewhite:2002:NWT

[App02d] Ashton Applewhite. News: WCDMA and the 3G turf wars. IEEE Distributed Systems Online, 3(12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0212/f/news_print.htm.

Applewhite:2003:NSP

[App03] Ashton Applewhite. News: From T-shirts to pinstripes: P2P gets some respect. IEEE Distributed Systems Online, 4 (1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0301/f/news_print.htm.

Assad:2007:DMM

[AR07] André Elia Assad and Nelson Souto Rosa. Device middleware for mobile applications. IEEE Distributed Systems Online, 8(8):??, August 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.

org/comp/mags/ds/2007/08/
mds2007080001.pdf.

Arnedillo-Sanchez:2007:ML

[AS07] Inmaculada Arnedillo-Sánchez.

Mobile learning. IEEE Distributed Systems Online, 8(6): ??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/06/o6001.pdf.

Ashrafi:2004:OOD

[ATS04] M. Z. Ashrafi, D. Taniar, and K. Smith. ODAM: An optimized distributed association rule mining algorithm. IEEE Distributed Systems Online, 5 (3):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28687/01285877.pdf?isnumber=
28687&prod=JNL&arnumber=1285877&
arSt=&ared=&arAuthor=Ashrafi%
2C+M.Z.%3B+Taniar%2C+D.%3B+
Smith%2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28687&arnumber=1285877&
count=7&index=0;   http:/
/ieeexplore.ieee.org/xpls/
references.jsp?arnumber=1285877.

Attiya:2007:CPD

[Att07] Hagit Attiya. Concurrency and the principle of data locality. IEEE Distributed Systems Online, 8(9):??, September 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.

computer.org/comp/mags/ds/
2007/09/mds2007090003.pdf.

Almeida:2005:DIS

[AvSQP05] J. P. Almeida, M. van Sinderen, D. A. C. Quartel, and L. F. Pires. Designing interaction systems for distributed applications. IEEE Distributed Systems Online, 6(3):??, March 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30740/01423916.pdf?isnumber=
30740&prod=JNL&arnumber=1423916&
arSt=&ared=&arAuthor=Almeida%
2C+J.P.%3B+van+Sinderen%2C+
M.%3B+Quartel%2C+D.A.C.%3B+
Pires%2C+L.F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30740&arnumber=1423916&
count=4&index=0.

Ben-Ari:2001:EI

[BAB01] Mordechai Ben-Ari and Alan Burns. Extreme interleavings. IEEE Distributed Systems Online, 2(1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0101/highlights/edu0101.
htm.

Bacon:2000:I

[Bac00]  Jean Bacon.  Introduction.

IEEE Distributed Systems Online, 1(1), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds100/ds1eicprint.
htm.

Bacon:2002:DON

[Bac02]  Jean Bacon.  DS Online’s

new partners. IEEE Distributed Systems Online, 3 (1), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0201/features/eic_1.htm.

Bacon:2004:MWP

[Bac04] J. Bacon. Middleware 2004 works in progress. IEEE Distributed Systems Online, 5(8): 1, August 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29550/01341359.pdf?isnumber=
29550&prod=JNL&arnumber=1341359&
arSt=+1&ared=+1&arAuthor=+
Bacon%2C+J.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29550&arnumber=1341359&
count=5&index=0.

Badica:2007:MCP

[Bad07] Costin Badica. Manufacturing and control: Putting agents to work. IEEE Distributed Systems Online, 8(6):??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/06/o6005.pdf.

Buyya:2001:NGR

[BAG01] Rajkumar Buyya, David Abramson, and Jonathan Giddy. Nimrod-G resource broker for service-oriented Grid computing. IEEE Distributed Systems Online, 2(7), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0107/departments/res0107_
print.htm.

Baker:2001:RDS

[Bak01] Mark Baker. Review: Distributed systems textbook falls short of its full potential. IEEE Distributed Systems Online, 2(4), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0104/departments/bks0104a.
htm.

Baker:2004:IFR

[Bak04] M. Baker. Ian Foster on recent changes in the Grid community. IEEE Distributed Systems Online, 5(2):4/1–4/10, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28454/01270784.pdf?isnumber=
28454&prod=JNL&arnumber=1270784&
arSt=+4%2F1&ared=+4%2F10&
arAuthor=+Baker%2C+M.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270784&count=6&
index=2.

Banavar:2003:WPMa

[Ban03a] Guruduth Banavar. Works in progress: Middleware works in progress papers. IEEE Distributed Systems Online, 4 (5), 2003. CODEN ????

ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/WIPs/index_may.htm.

Banavar:2003:WPMb

[Ban03b] Guruduth Banavar. Works in progress: Middleware works in progress papers. IEEE Distributed Systems Online, 4 (6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/WIPs/index.htm.

Barake:2003:BRE

[Bar03] Tony Barake. Book review: The enterprise goes to school: a review of Wireless Java Programming for Enterprise Applications: Mobile Devices Go Corporate by Dan Harkey, Shan Appajodu, and Mike Larkin. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/d/bks_a.htm.

Bertino:2002:DTW

[BB02] Elisa Bertino and Athman Bouguettaya. Database technology on the Web. IEEE Distributed Systems Online, 3 (7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/departments/w4gei.
htm.

Bellotti:2002:UTH

[BBdGM02] Francesco Bellotti, Riccardo Berta, Alessandro de Gloria, and Massimiliano Margarone. User testing a hypermedia tour guide. IEEE Distributed Systems Online, 3 (6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0206/features/bp2bel.
htm.

Blair:2001:DIO

[BCA^{+}01] Gordon S. Blair, Geoff Coulson, Anders Andersen, Lynne Blair, Michael Clarke, Fabio Costa, Hector Duran-Limon, Tom Fitzpatrick, Lee Johnston, Rui Moreira, Nikos Parlavantzas, and Katia Saikoski. The design and implementation of Open ORB 2. IEEE Distributed Systems Online, 2 (6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0106/features/bla0106_
print.htm.

Blair:2001:RMa

[BCCK01a] Gordon Blair, Fabio Costa, Roy Campbell, and Fabio Kon. Reflective middleware. IEEE Distributed Systems Online, 2(5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/gei0105.
htm.

Blair:2001:RMb

[BCCK01b] Gordon Blair, Fabio Costa, Roy Campbell, and Fabio Kon. Reflective middleware. IEEE Distributed Systems Online, 2(6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/gei0105.
htm.

Billinghurst:2003:ICG

[BCPK03] Mark Billinghurst, Adrian Cheok, Simon Prince, and Hirokazu Kato. IEEE computer graphics & applications: Projects in VR: Real world teleconferencing. IEEE Distributed Systems Online, 4 (2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/d/bill_print.htm.

Bellavista:2002:UPI

[BCS02] Paolo Bellavista, Antonio Corradi, and Cesare Stefanelli. The ubiquitous provisioning of Internet services to portable devices. IEEE Distributed Systems Online, 3(9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/f/bp3bel.htm.

Berzal:2004:STC

[Ber04]  F. Berzal.  Standards for

Talking Computers [book review]. IEEE Distributed Systems Online, 5(3):1–3, ????

2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28687/01285882.pdf?isnumber=
28687&prod=JNL&arnumber=1285882&
arSt=+1&ared=+3&arAuthor=
Berzal%2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28687&arnumber=1285882&
count=7&index=5.

Berzal:2005:EW

[Ber05a]  F. Berzal.  The editable

Web. IEEE Distributed Systems Online, 6(3):6, March 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30740/01423921.pdf?isnumber=
30740&prod=JNL&arnumber=1423921&
arSt=+6&ared=+6&arAuthor=
Berzal%2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30740&arnumber=1423921&
count=4&index=3.

Berzal:2005:JTF

[Ber05b] F. Berzal. Java tools and frameworks for Web application development. IEEE Distributed Systems Online, 6(5): 5, May 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/31212/01453474.pdf?isnumber=
31212&prod=JNL&arnumber=1453474&
arSt=+5&ared=+5&arAuthor=+
Berzal%2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31212&arnumber=1453474&
count=4&index=2.

Bao:2005:HSU

[BFHX05] Xiuguo Bao, Binxing Fang, Mingzhen Hu, and Binbin Xu. Heterogeneous search in unstructured peer-to-peer networks. IEEE Distributed Systems Online, 6(2):??, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30522/01407758.pdf?isnumber=
30522&prod=JNL&arnumber=1407758&
arSt=&ared=&arAuthor=Xiuguo+
Bao%3B+Binxing+Fang%3B+Mingzhen+
Hu%3B+Binbin+Xu;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30522&
arnumber=1407758&count=5&
index=0.

Baliga:2004:SCN

[BGSK04] G. Baliga, S. Graham, Lui Sha, and P. R. Kumar. Service continuity in networked control using Etherware. IEEE Distributed Systems Online, 5(9):2, September 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29771/01355902.pdf?isnumber=
29771&prod=JNL&arnumber=1355902&
arSt=+2&ared=+2&arAuthor=
+Baliga%2C+G.%3B++Graham%
2C+S.%3B++Lui+Sha%3B++Kumar%
2C+P.R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29771&arnumber=1355902&
count=4&index=1;   http:/
/ieeexplore.ieee.org/xpls/
references.jsp?arnumber=1355902.

Bing:2004:GWT

[Bin04] B. Bing. A good wireless technologies reference. IEEE Distributed Systems Online, 5(5): 6, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/28996/01306408.pdf?isnumber=
                                        28996&prod=JNL&arnumber=1306408&
                                        arSt=+6&ared=+6&arAuthor=
                                        Bing%2C+B.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=28996&arnumber=1306408&
                                        count=5&index=4.

Birman:2003:IIC

[Bir03] Ken Birman. IEEE Internet computing: Peer to peer: The League of SuperNets. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/f/w5p2pp.htm.

Buyya:2000:RPP

[BJ00] Rajkumar Buyya and Hai Jin. Review: Parallel Programming: Techniques and Applications Using Networked Workstations and Parallel Computers. IEEE Distributed Systems Online, 1(3), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds300/ds3book.
                                        htm.

Berzal:2001:TTJ

[BM01] Fernando Berzal and Nicolás Marín. Tools and tips for Java programming. IEEE Distributed Systems Online, 2 (5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/departments/bks0105a.
htm.

Bratus:2007:HCHb

[BM07] Sergey Bratus and Christopher Masone. Hacker curriculum: How we can use it in teaching. IEEE Distributed Systems Online, 8(11):??, November 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/11/mds2007110002.pdf.

Bond:2003:BRC

[Bon03]  Tim Bond.  Book review:

Cooking with XSLT: a review of XSLT Cookbook by Sal Mangano. IEEE Distributed Systems Online, 4 (9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/0306/d/bks_a.htm.

Bond:2006:BRC

[Bon06] Tim Bond. Book reviews:

Component-oriented programming: a .NET perspective. IEEE Distributed Systems Online, 7(5):??, May 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.

                                        org/comp/mags/ds/2006/05/
                                        o5006.pdf.

Buyya:2006:CPC

[BPBT06] Rajkumar Buyya, Al-Mukaddim Khan Pathan, James Broberg, and Zahir Tari. A case for peering of content delivery networks. IEEE Distributed Systems Online, 7(10):??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/10/ox003.pdf.

Baldoni:2002:FDC

[BR02] Roberto Baldoni and Michel Raynal. Fundamentals of distributed computing: a practical tour of vector clock systems. IEEE Distributed Systems Online, 3(2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0202/features/bal_print.
                                        htm.

Buyya:2007:IMP

[BR07] Rajkumar Buyya and Kotagiri Ramamohanarao. An innovative master’s program in distributed computing. IEEE Distributed Systems Online, 8(1): ??, January 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2007/01/o1002.pdf.

Bratus:2007:HCHa

[Bra07] Sergey Bratus. Hacker curriculum : How hackers learn networking. IEEE Distributed Systems Online, 8(10):??, October 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/10/mds2007100002.pdf.

Byrne:2007:ML

[BTSV^{+}07] Peter Byrne, Brendan Tangney, Pedro P. Sanchez-Villalon, Manuel Ortega, Lin-Jung Wu, Kao-En Chang, Hsien-Sheng Hsiao, and Yao-Ting Sung. Mobile learning. IEEE Distributed Systems Online, 8(7):??, July 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/07/mds2007070003.pdf.

Buyya:2001:VLP

[Buy01] Rajkumar Buyya. The Virtual Laboratory Project. IEEE Distributed Systems Online, 2 (5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/tvl0105.
htm.

Buyya:2007:CGC

[BV07] Rajkumar Buyya and Srikumar Venugopal. Cluster and Grid computing: a graduate distributed-computing course. IEEE Distributed Systems Online, 8(12):??, December 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.

                                        computer.org/comp/mags/ds/
                                        2007/12/mds2007120002.pdf.

Curbera:2002:UWS

[C^{+}02] Francisco Curbera et al. Unraveling the Web services web: An introduction to SOAP, WSDL, and UDDI. IEEE Distributed Systems Online, 3(4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0204/features/wp2spot.
                                        htm.

C:2003:BRH

[C.03] Héctor Zenil C. Book review: How to secure your enterprise: a review of Securing Business Information: Strategies to Protect the Enterprise and Its Network by F. Christian Byrnes and Dale Kutnick. IEEE Distributed Systems Online, 4(3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0303/d/bks_b.htm.

Cai:2003:BRD

[Cai03] Min Cai. Book review: Distributed computing basics: a review of Distributed Computing: Principles and Applications by M. L. Liu. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/d/bks_a.htm.

Caprio:2008:PM

[Cap08] Griffin Caprio. Parallel metaheuristics. IEEE Distributed Systems Online, 9(8):??, August 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Chandy:2002:UCF

[CFJ^{+}02] Mani Chandy, Axel Fuchs, Bill Janssen, Deepak Mulchandani, and Mark Weiser. Ubiquitous computing: The future of development? IEEE Distributed Systems Online, 3 (3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/ic/wp3round.htm.

Chavez:2003:BRH

[Cha03] Hector Zenil Chavez. Book review: How to develop enterprise-secure Java applications learning risks A review of Hacking Exposed J2EE and Java, Developing Secure Applications with Java Technology by Art Taylor, Brian Buege, and Randy Layman. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/d/bks_c.htm.

Chavez:2004:APC

[Cha04a] H. Z. Chavez. Applied Parallel Computing [book review]. IEEE Distributed Systems Online, 5(3):1–3, ????

2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28687/01285883.pdf?isnumber=
28687&prod=JNL&arnumber=1285883&
arSt=+1&ared=+3&arAuthor=
Chavez%2C+H.Z.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28687&
arnumber=1285883&count=7&
index=6.

Chavez:2004:AIP

[Cha04b] H. Z. Chavez. Artificial intelligence profits from biological lessons. IEEE Distributed Systems Online, 5(5):5, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28996/01306407.pdf?isnumber=
28996&prod=JNL&arnumber=1306407&
arSt=+5&ared=+5&arAuthor=
Chavez%2C+H.Z.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28996&
arnumber=1306407&count=5&
index=3.

Chavez:2004:GKY

[Cha04c] H. Z. Chavez. Getting to know your enemy. IEEE Distributed Systems Online, 5(7): 6, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29290/01323041.pdf?isnumber=
29290&prod=JNL&arnumber=1323041&
arSt=+6&ared=+6&arAuthor=
Chavez%2C+H.Z.;   http://
ieeexplore.ieee.org/xpls/

abs_all.jsp?isnumber=29290&
arnumber=1323041&count=4&
index=3.

Chavez:2004:SF

[Cha04d] H. Z. Chavez. Security fundamentals. IEEE Distributed Systems Online, 5(6):3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/29028/01308175.pdf?isnumber=
29028&prod=JNL&arnumber=1308175&
arSt=+3&ared=+3&arAuthor=
Chavez%2C+H.Z.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29028&
arnumber=1308175&count=2&
index=0.

Chavez:2005:JFE

[Cha05]  H. Z. Chavez.  JavaServer

Faces: An excellent learning tool. IEEE Distributed Systems Online, 6(2):6, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30522/01407763.pdf?isnumber=
30522&prod=JNL&arnumber=1407763&
arSt=+6&ared=+6&arAuthor=
Chavez%2C+H.Z.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30522&
arnumber=1407763&count=5&
index=4.

Chen:2002:JCN

[Che02]  Jiadong Chen.  Java E-

commerce in a nutshell: a review of Professional Java E-Commerce. IEEE Distributed Systems Online, 3 (4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0204/departments/bks_
                                        a.htm.

Che:2005:CPT

[Che05a] Haoyang Che. A cornucopia of Perl techniques. IEEE Distributed Systems Online, 6(1): 6, January 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/30521/01407757.pdf?isnumber=
                                        30521&prod=JNL&arnumber=1407757&
                                        arSt=+6&ared=+6&arAuthor=
                                        Haoyang+Che;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30521&arnumber=1407757&
                                        count=4&index=3.

Che:2005:MDC

[Che05b] Haoyang Che. Mastering distributed computing. IEEE Distributed Systems Online, 6(3): 5, March 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/30740/01423920.pdf?isnumber=
                                        30740&prod=JNL&arnumber=1423920&
                                        arSt=+5&ared=+5&arAuthor=
                                        Haoyang+Che;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30740&arnumber=1423920&
                                        count=4&index=2.

Coulson:2006:SNG

[CKB06] Geoff Coulson, Dean Kuo, and John Brooke. Sensor networks + Grid computing = A new challenge for the Grid? IEEE Distributed Systems Online, 7 (12):??, December 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/12/
oz002.pdf.

Constantine:2002:UCE

[CL02] Larry L. Constantine and Lucy A. D. Lockwood. Usagecentered engineering for Web applications. IEEE Distributed Systems Online, 3 (3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/features/con_print.
htm.

Clark:2003:IIC

[Cla03] David D. Clark. IEEE Internet computing: Peer to peer: Economics and the design of open systems. IEEE Distributed Systems Online, 4 (4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/f/p2p_print.htm.

Curry:2006:WPI

[CM06] Edward Curry and Jacques Mossière. Works in progress: The 2nd International Middleware Doctoral Symposium. IEEE Distributed Systems Online, 7(3):??, March 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/03/
o3004.pdf.

Cole:2004:ULS

[Col04] B. Cole. Utilities look to the skies for monitoring the power grid. IEEE Distributed Systems Online, 5(11):3, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30052/01377090.pdf?isnumber=
30052&prod=JNL&arnumber=1377090&
arSt=+3&ared=+3&arAuthor=+
Cole%2C+B.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30052&arnumber=1377090&
count=5&index=2.

Cortes:2004:BIL

[Cor04] J. A. Cortes. A brief introduction to LDAP and its future. IEEE Distributed Systems Online, 5(4):1–3, April 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28913/01301257.pdf?isnumber=
28913&prod=JNL&arnumber=1301257&
arSt=+1&ared=+3&arAuthor=
Cortes%2C+J.A.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28913&
arnumber=1301257&count=5&
index=4.

Costlow:2002:BSC

[Cos02]  Terry Costlow.  In brief:

Spam comes to mobile phones. IEEE Distributed Systems Online, 3(12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0212/d/brief1.htm.

Costlow:2003:BHG

[Cos03a] Terry Costlow. In brief: Hotels go wireless. IEEE Distributed Systems Online, 4 (3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0303/d/brief.htm.

Costlow:2003:BIP

[Cos03b] Terry Costlow. In brief: Idle PCs might be key to smallpox cure. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/d/brief.htm.

Costlow:2003:BIM

[Cos03c] Terry Costlow. In brief: Intelligent mail idea raising opponents’ ire. IEEE Distributed Systems Online, 4 (9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/d/brief.htm.

Costlow:2003:BPA

[Cos03d] Terry Costlow. In brief: Piracy, advertising groups target spyware. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0306/d/brief.htm.

Costlow:2003:BPL

[Cos03e] Terry Costlow. In brief: Portability lets telemarketers target cell phones. IEEE Distributed Systems Online, 4 (11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/d/brief.htm.

Costlow:2004:CPB

[Cos04a] T. Costlow. Camera phone bans expected. IEEE Distributed Systems Online, 5(2):5/1–5/3, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28454/01270785.pdf?isnumber=
28454&prod=JNL&arnumber=1270785&
arSt=+5%2F1&ared=+5%2F3&arAuthor=
+Costlow%2C+T.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270785&count=6&
index=3.

Costlow:2004:DPH

[Cos04b]  T. Costlow.  Digital post-

mark helps fight spam, virus attacks. IEEE Distributed Systems Online, 5(5):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28996/01306406.pdf?isnumber=
28996&prod=JNL&arnumber=1306406&
arSt=&ared=&arAuthor=Costlow%

2C+T.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28996&arnumber=1306406&
count=5&index=2.

Costlow:2004:VTB

[Cos04c]  T. Costlow.  Virtual table

brings distant loved ones together. IEEE Distributed Systems Online, 5(1):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28452/01270715.pdf?isnumber=
28452&prod=JNL&arnumber=1270715&
arSt=&ared=&arAuthor=Costlow%
2C+T.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28452&arnumber=1270715&
count=8&index=4.

Costlow:2004:WFC

[Cos04d] T. Costlow. Wi-fi could augment GPS. IEEE Distributed Systems Online, 5(9):3, September 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/29771/01355903.pdf?isnumber=
29771&prod=JNL&arnumber=1355903&
arSt=+3&ared=+3&arAuthor=
+Costlow%2C+T.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29771&
arnumber=1355903&count=4&
index=2.

Coulouris:2000:SEP

[Cou00]  George Coulouris.  Sys-

tems for electronic publishing. IEEE Distributed Systems Online, 1(1), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds100/ds1opiprint.
                                        htm.

Chen:2003:IIC

[CP03] Yih-Farn Robin Chen and Charles Petrie. IEEE Internet computing: Guest Editors’ introduction: Ubiquitous mobile computing. IEEE Distributed Systems Online, 4 (4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0304/f/w2gei.htm.

Chrysanthis:2003:GEI

[CSZ03] Panos K. Chrysanthis, Morris Sloman, and Arkady Zaslavsky. Guest Editors’ introduction: Mobile data management. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/f/gei.htm.

Curran:2007:GM

[Cur07] Kevin Curran. Geographical mashups. IEEE Distributed Systems Online, 8(4):??, April 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/04/o4005.pdf.

Davids:2003:BSD

[Dav03] Angela Davids. In brief: Spam debate simmers during Congressional recess. IEEE Distributed Systems Online, 4 (8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/0308/d/brief2.htm.

Dave:2004:NPN

[Dav04a] M. A. Dave. Numerical programming on networks of computers. IEEE Distributed Systems Online, 5(11):4, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30052/01377091.pdf?isnumber=
30052&prod=JNL&arnumber=1377091&
arSt=+4&ared=+4&arAuthor=+
Dave%2C+M.A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30052&arnumber=1377091&
count=5&index=3.

Davids:2004:FOI

[Dav04b] A. Davids. Fraudulent Online Identity Sanctions Act: empowering law enforcement or limiting privacy? IEEE Distributed Systems Online, 5(4): ??, April 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28913/01301255.pdf?isnumber=
28913&prod=JNL&arnumber=1301255&
arSt=&ared=&arAuthor=Davids%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28913&arnumber=1301255&
count=5&index=2.

Dragoi:2004:DSE

[DB04] O. A. Dragoi and J. P. Black. Discovering services is not enough. IEEE Distributed Systems Online, 5(8):3, August 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/29550/01341361.pdf?isnumber=
                                        29550&prod=JNL&arnumber=1341361&
                                        arSt=+3&ared=+3&arAuthor=
                                        +Dragoi%2C+O.A.%3B++Black%
                                        2C+J.P.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29550&arnumber=1341361&
                                        count=5&index=2;   http:/
                                        /ieeexplore.ieee.org/xpls/
                                        references.jsp?arnumber=1341361.

deCamargo:2006:SCS

[dCKC06] Raphael Y. de Camargo, Fabio Kon, and Renato Cerqueira. Strategies for checkpoint storage on opportunistic Grids. IEEE Distributed Systems Online, 7(9):??, September 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/09/o9001.pdf.

Douglis:2003:IIC

[DF03] Fred Douglis and Ian Foster. IEEE Internet computing: Guest Editors’ introduction: The Grid grows up. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/0307/f/w4gei.htm.

Denny:2003:MDM

[DFCP03] Matthew Denny, Michael J.

Franklin, Paul Castro, and Apratim Purakayastha. Mobile data management (part 2): The design of a scalable discovery service for moving networked resources. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/f/denp.htm.

Drabkin:2008:RTC

[DFK08] V. Drabkin, R. Friedman, and G. Kliot. On the (un)reliability of TCP connections: The return of the end-to-end argument. IEEE Distributed Systems Online, 9(8):??, August 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Duran-Limon:2004:ARM

[DLBC04] H. A. Duran-Limon, G. S. Blair, and G. Coulson. Adaptive resource management in middleware: a survey. IEEE Distributed Systems Online, 5(7): ??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29290/01323036.pdf?isnumber=
29290&prod=JNL&arnumber=1323036&
arSt=&ared=&arAuthor=Duran-
Limon%2C+H.A.%3B+Blair%2C+

G.S.%3B+Coulson%2C+G.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29290&
arnumber=1323036&count=4&
index=0;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1323036.

Doswell:2002:IPC

[DMJ^{+}02] Jayfus Tucker Doswell, Tom Martin, Mark Jones, David A. Ross, Bruce B. Blasch, Daniel Siewiorek, and Asim Smailagic. IEEE pervasive computing: Wearable computing: Works in progress: Augmented-reality systems for treating drug dependency; designing E-textile architectures for wearable computing; integrated wearable computer helps users find their way; DiMA: Diabetes Management Assistant wearable computer. IEEE Distributed Systems Online, 3 (12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0212/f/b4wips.htm.

Denis:2003:PPP

[DPPR03] Alexandre Denis, Christian Pérez, Thierry Priol, and André Ribes. Parallel programming: Parallel CORBA objects for programming computational Grids. IEEE Distributed Systems Online, 4 (2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/f/pri_print.htm.

daRocha:2006:MCM

[dRE06] Ricardo Couto Antunes da Rocha and Markus Endler. Middleware: Context management in heterogeneous, evolving ubiquitous environments. IEEE Distributed Systems Online, 7(4): ??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/04/o4001.pdf.

Drusinsky:2007:CVE

[DSD07] Doron Drusinsky, Man-Tak Shing, and Kadir Alpaslan Demir. Creating and validating embedded assertion statecharts. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/05/o5003.pdf.

Detmold:2008:MDV

[DvdHD^{+}08] Henry Detmold, Anton van den Hengel, Anthony Dick, Katrina Falkner, David S. Munro, and Ron Morrison. Middleware for distributed video surveillance. IEEE Distributed Systems Online, 9(2):??, February 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/02/mds2008020001.pdf.

Dwivedi:2001:RCG

[Dwi01] Neelam Dwivedi. Review: A Comprehensive Guide for Beginners in WAP. IEEE Distributed Systems Online, 2 (3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0103/departments/bks0103.
                                        htm.

Dwivedi:2002:BRB

[Dwi02a] Neelam Dwivedi. Book review: Brilliant book on rules of ownership: a review of Cyber Regs: A Business Guide to Web Property, Privacy and Patents by Bill Zoellick. IEEE Distributed Systems Online, 3(12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0212/d/bks_a.htm.

Dwivedi:2002:BRO

[Dwi02b] Neelam Dwivedi. Book review: One-stop reference for networking fundamentals: a review of Computer Networks and Internets with Internet Applications, 3rd ed. by Douglas E. Comer. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0210/d/bks_b.htm.

El-Alfy:2007:RNS

[EA07] El-Sayed M. El-Alfy. A review of network security. IEEE Distributed Systems Online, 8(7):??, July 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/07/
mds2007070006.pdf.

Epp:2003:IPC

[Epp03] Edward C. Epp. IEEE pervasive computing: Identity management: Relationship management: Secure collaboration in a ubiquitous environment. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/f/epp_print.htm.

Estevez-Tapiador:2003:BRC

[ET03] Juan M. Estevez-Tapiador.

Book review: Concepts and attitudes for Internet security: a review of Firewalls and Internet Security: Repelling the Wily Hacker, second edition by William R. Cheswick, Steven M. Bellovin, and Aviel D. Rubin. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0310/d/bks_b.htm.

Estevez-Tapiador:2004:ECT

[ET04a] J. M. Estevez-Tapiador. The emergence of cyber-terrorism. IEEE Distributed Systems Online, 5(10):4, October 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/iel5/8968/29772/01355909.
                                        pdf?isnumber=29772&prod=JNL&
                                        arnumber=1355909&arSt=+4&
                                        ared=+4&arAuthor=+Estevez-
                                        Tapiador%2C+J.M.;   http:/
                                        /ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29772&
                                        arnumber=1355909&count=3&
                                        index=2.

Estevez-Tapiador:2004:WSS

[ET04b] J. M. Estevez-Tapiador. Web Services Security [book review]. IEEE Distributed Systems Online, 5(1):8.1–8.4, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/28452/01270717.pdf?isnumber=
                                        28452&prod=JNL&arnumber=1270717&
                                        arSt=+8.1&ared=+8.4&arAuthor=
                                        Estevez-Tapiador%2C+J.M.;
                                         http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        28452&arnumber=1270717&count=
                                        8&index=6.

Estevez-Tapiador:2006:BRP

[ET06] Juan M. Estevez-Tapiador. Book review: Practical intrusion prevention. IEEE Distributed Systems Online, 7(6): ??, June 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/06/o6005.pdf.

Farber:2002:BSL

[Far02] David Farber. Balancing security and liberty. IEEE Distributed Systems Online, 3 (2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0202/departments/wp102ic.
htm.

Fergusson:2008:DCEd

[FBG^{+}08] D. Fergusson, R. Barbera, E. Giorgio, M. Fargetta, G. Sipos, D. Romano, M. Atkinson, and E. V. Meer. Distributed computing education, part 4: Training infrastructure. IEEE Distributed Systems Online, 9(10):??, October 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Flores-Cortes:2007:AMO

[FCBG07] Carlos A. Flores-Cortés, Gordon S. Blair, and Paul Grace. An adaptive middleware to overcome service discovery heterogeneity in mobile ad hoc environments. IEEE Distributed Systems Online, 8(7):??, July 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/07/mds2007070001.pdf.

Feitelson:2007:TTI

[Fei07] Dror Feitelson. Teaching TCP/ IP hands-on. IEEE Distributed Systems Online, 8(11): ??, November 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/11/
mds2007110005.pdf.

Ferrero:2001:NMH

[Fer01]  Jenny Ferrero.  No more

hanging chads. IEEE Distributed Systems Online, 2 (3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0103/features/vot0103.
                                        htm.

Feustel:2002:WSJ

[Feu02] Edward Feustel. Writing secure Java applications: a review of Java Security, 2nd ed. IEEE Distributed Systems Online, 3(3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0203/departments/bks_
                                        a.htm.

Farkas:2003:IPC

[FH03] Keith Farkas and Guerney Hunt. IEEE pervasive computing: New products. IEEE Distributed Systems Online, 4 (10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/d/bpro.htm.

Friedman:2006:DWA

[FH06] Roy Friedman and Erez Hadad. Distributed wisdom: Analyzing distributed-system performance — latency vs. throughput. IEEE Distributed Systems Online, 7(1):??, January 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/01/o1001.pdf.

Fergusson:2008:DCEb

[FHR^{+}08] D. Fergusson, R. Hopkins, D. Romano, E. Vander Meer, and M. Atkinson. Distributed computing education, part 2: International summer schools. IEEE Distributed Systems Online, 9(7):??, July 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Fewings:2007:DGP

[FJ07] Ade J. Fewings and Nigel W.

John. Distributed graphics pipelines on the Grid. IEEE Distributed Systems Online, 8 (1):??, January 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/01/
o1001.pdf.

Fergusson:2008:DCEc

[FJH^{+}08] D. Fergusson, P. Jandric, R. Hopkins, E. Vander Meer, and M. Atkinson. Distributed computing education, part 3: The winter school online experience. IEEE Distributed Systems Online, 9(9):??, September 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Fox:2001:IWC

[Fox01] Jeffrey Fox. Introduction to Web computing. IEEE Distributed Systems Online, 2 (2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0102/highlights/web0102.
htm.

Fox:2002:MPP

[Fox02] Geoffrey Fox. Message passing: From parallel computing to the Grid. IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0208/d/cp5web.htm.

Fox:2003:CSE

[Fox03] Geoffrey Fox. Computing in science & engineering: Web computing: Java and Grande applications. IEEE Distributed Systems Online, 4 (4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/d/web_print.htm.

Frei:2005:EAA

[FPA05] A. Frei, A. Popovici, and G. Alonso. Eventizing applications in an adaptive middleware platform. IEEE Distributed Systems Online, 6(4): 1, April 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/30778/01425404.pdf?isnumber=
30778&prod=JNL&arnumber=1425404&
arSt=+1&ared=+1&arAuthor=
+Frei%2C+A.%3B++Popovici%
2C+A.%3B++Alonso%2C+G.;

http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
30778&arnumber=1425404&count=
3&index=0.

Francis:2003:IIC

[Fra03] Paul Francis. IEEE Internet computing: Peer to peer: Is the Internet going NUTSS? IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/f/w6p2pp.htm.

Fergusson:2008:DCEa

[FvdMAR08] David Fergusson, Elizabeth van der Meer, Malcolm Atkinson, and Diego Romano. Distributed computing education, part 1: a special case? IEEE Distributed Systems Online, 9(6):??, June 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/06/
mds2008060002.pdf.

Galant:2005:BCT

[Gal05]  V. Galant.  Blending E-

commerce theory and application. IEEE Distributed Systems Online, 6(1):5, January 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30521/01407756.pdf?isnumber=
30521&prod=JNL&arnumber=1407756&
arSt=+5&ared=+5&arAuthor=
Galant%2C+V.;   http://ieeexplore.

                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30521&arnumber=1407756&
                                        count=4&index=2.

Gama:2006:SLV

[Gam06] Pedro Gama. Simplifying largevolume data sharing in distributed environments. IEEE Distributed Systems Online, 7 (1):??, January 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/01/
                                        o1005.pdf.

Gama:2007:SRW

[Gam07] Pedro Gama. Stop reinventing the wheel. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/05/o5009.pdf.

Ganzha:2006:BRW

[Gan06] Maria Ganzha. Book reviews: When humans and agents collide. IEEE Distributed Systems Online, 7(3):??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/03/o3009.pdf.

Ganzha:2007:BRS

[Gan07] Maria Ganzha. Book review: Software agent technology: The state of the art. IEEE Distributed Systems Online, 8(1): ??, January 2007. CODEN ????

ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/01/o1005.pdf.

Ganzha:2008:LIA

[Gan08]  Maria Ganzha.  E-learning

with intelligent agents. IEEE Distributed Systems Online, 9 (2):??, February 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/02/
mds2008020004.pdf.

Gavrila:2001:SBP

[Gav01] Dariu M. Gavrila. Sensorbased pedestrian protection. IEEE Distributed Systems Online, 2(8), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0108/departments/its108.
htm.

Gartner:2003:DSP

[GBK03] Felix C. Gärtner, Levente Buttyán, and Klaus Kursawe. Dependable systems: PoDSy workshop report: From fault tolerance to security and back. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/f/gar_intro.htm.

Geer:2006:BGB

[Gee06] David Geer. In brief: Gridbased modeling technique furthers earthquake risk prediction. IEEE Distributed Systems Online, 7(2):??, February 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/02/o2005.pdf.

Gentzsch:2002:GCV

[Gen02] Wolfgang Gentzsch. Grid computing: a vendor’s vision. IEEE Distributed Systems Online, 3(6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0206/departments/new.
htm.

Gawinecki:2008:MAS

[GF08] Maciej Gawinecki and Grzegorz Frackowiak. Multi-agent systems with JADE: a guide with extensive study. IEEE Distributed Systems Online, 9 (3):??, March 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/03/
mds2008030004.pdf.

Ghose:2003:MDM

[GGC03] Abhishek Ghose, Jens Grossklags, and John Chuang. Mobile data management (part 2): Resilient data-centric storage in wireless sensor networks. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/f/jenp.htm.

Gilreath:2000:RDP

[Gil00]  William Gilreath.  Review:

Distributed Programming with Java. IEEE Distributed Systems Online, 1(1), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds100/ds1bookprint.
htm.

Gilreath:2001:JNP

[Gil01] William F. Gilreath. Java Network Programming introduces the fundamentals. IEEE Distributed Systems Online, 2(5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/departments/bks0105b.
htm.

Gerber:2002:ABI

[GK02] Andreas Gerber and Matthias Klusch. Agent-based integrated services for timber production and sales. IEEE Distributed Systems Online, 3 (2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0202/features/ger_print.
htm.

Graf:2007:GIE

[GMG07] Philipp Graf and Klaus D.

Müller-Glaser. Gaining insight into executable models during runtime: Architecture and mappings. IEEE Distributed Systems Online, 8(3):??, March 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/03/o3001.pdf.

Gong:2002:PPN

[Gon02] Li Gong. Peer to peer networks in action. IEEE Distributed Systems Online, 3 (1), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0201/ic/w102gei.htm.

Gordon:2004:IRT

[Gor04a] M. Gordon. An introduction to RDF technologies: Too little, too soon. IEEE Distributed Systems Online, 5(8): 5, August 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29550/01341363.pdf?isnumber=
29550&prod=JNL&arnumber=1341363&
arSt=+5&ared=+5&arAuthor=+
Gordon%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29550&arnumber=1341363&
count=5&index=4.

Gordon:2004:KRL

[Gor04b] M. Gordon. Knowledge Representation: Logical, Philosophical, and Computational Foundations [book review]. IEEE Distributed Systems Online, 5 (1):9.1–9.3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/28452/01270718.
pdf?isnumber=28452&prod=JNL&
arnumber=1270718&arSt=+9.
1&ared=+9.3&arAuthor=Gordon%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28452&arnumber=1270718&
count=8&index=7.

Gordon:2007:TK

[Gor07] Minor Gordon. Twisted in knots. IEEE Distributed Systems Online, 8(3):??, March 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/03/o3005.pdf.

Goth:2002:AO

[Got02a]  Greg Goth.  Are the air-

waves overcrowded? IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/f/news_print.htm.

Goth:2002:GSM

[Got02b] Greg Goth. GovNet is still marking time. IEEE Distributed Systems Online, 3 (4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/departments/new.
htm.

Goth:2002:NER

[Got02c]  Greg Goth.  A new era

for research and development? IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0205/departments/new_
                                        print.htm.

Goth:2002:NMB

[Got02d] Greg Goth. News: Mobile backlash. IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0211/f/news_print.htm.

Goth:2002:WSS

[Got02e] Greg Goth. Wireless security still ad hoc and add-on. IEEE Distributed Systems Online, 3(7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0207/features/news_print.
                                        htm.

Goth:2003:NOS

[Got03a] Greg Goth. News: Oh, say should you G? IEEE Distributed Systems Online, 4 (3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0303/f/news_print.htm.

Goth:2003:NWS

[Got03b] Greg Goth. News: Web services easing toward the mainstream. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0310/f/d10newp.htm.

Goth:2003:NWW

[Got03c] Greg Goth. News: Will the White House take on wi-fi? IEEE Distributed Systems Online, 4(2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/f/news_print.htm.

Goth:2004:CSW

[Got04a] G. Goth. Critics say Web services need a REST. IEEE Distributed Systems Online, 5(12): 1, December 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/30113/01381275.
pdf?isnumber=30113&prod=JNL&
arnumber=1381275&arSt=+1&
ared=+1&arAuthor=+Goth%2C+
G.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30113&arnumber=1381275&count=
4&index=0.

Goth:2004:WSM

[Got04b] G. Goth. When is a “Little in the Middle” ok? The Internet’s end-to-end principle faces more debate. IEEE Distributed Systems Online, 5(3):

??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28687/01285878.pdf?isnumber=
28687&prod=JNL&arnumber=1285878&
arSt=&ared=&arAuthor=Goth%
2C+G.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28687&arnumber=1285878&
count=7&index=1.

Goth:2004:WMS

[Got04c] G. Goth. Wireless MAN standard signals next-gen opportunities. IEEE Distributed Systems Online, 5(8):4, August 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/29550/01341362.pdf?isnumber=
29550&prod=JNL&arnumber=1341362&
arSt=+4&ared=+4&arAuthor=+
Goth%2C+G.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29550&arnumber=1341362&
count=5&index=3.

Goth:2005:OSM

[Got05] Greg Goth. Open source meets venture capital. IEEE Distributed Systems Online, 6(6): ??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2005/06/o6002.pdf.

Goth:2006:DTN

[Got06a] Greg Goth. Delay-tolerant network technologies coming together. IEEE Distributed Systems Online, 7(8):??, August 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/08/o8002.pdf.

Goth:2006:EPP

[Got06b] Greg Goth. From the Editors: Peer-to-peer community — looking beyond the legacy of Napster and Gnutella. IEEE Distributed Systems Online, 7 (3):??, March 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/03/
o3005.pdf.

Goth:2006:FMT

[Got06c] Greg Goth. Functionality meets terminology to address network security vulnerabilities. IEEE Distributed Systems Online, 7 (6):??, June 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/06/o6004.pdf.

Goth:2006:GIH

[Got06d] Greg Goth. Global information health networks get a boost. IEEE Distributed Systems Online, 7(7):??, July 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/07/
o7004.pdf.

Goth:2006:GCG

[Got06e] Greg Goth. Grid computing gets small. IEEE Distributed Systems Online, 7(11): ??, November 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/11/
                                        oy003.pdf.

Goth:2006:NCL

[Got06f] Greg Goth. News: Cradle of liberty lags on E-voting. IEEE Distributed Systems Online, 7(4):??, April 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/04/
                                        o4004.pdf.

Goth:2006:NDD

[Got06g] Greg Goth. News: Data-driven enterprise — slouching toward the Semantic Web. IEEE Distributed Systems Online, 7(3): ??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/03/o3006.pdf.

Goth:2006:NWD

[Got06h] Greg Goth. News: It’s a WLAN-derful life. IEEE Distributed Systems Online, 7(10): ??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/10/ox004.pdf.

Goth:2006:NNP

[Got06i] Greg Goth. News: New perils in hosted data? IEEE Distributed Systems Online, 7(5):??, May 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/05/o5004.pdf.

Goth:2006:NSN

[Got06j] Greg Goth. News: Not in the script — news of Java’s demise is premature. IEEE Distributed Systems Online, 7 (2):??, February 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/02/
o2004.pdf.

Goth:2006:NLS

[Got06k] Greg Goth. News: This little standard went to market; this little standard blew up. IEEE Distributed Systems Online, 7(12):??, December 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/12/oz005.pdf.

Goth:2006:NGL

[Got06l] Greg Goth. Next-gen landscape extends beyond the console. IEEE Distributed Systems Online, 7(9):??, September 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.

                                        computer.org/comp/mags/ds/
                                        2006/09/o9004.pdf.

Goth:2006:XCA

[Got06m] Greg Goth. XML: The center of attention up and down the stack. IEEE Distributed Systems Online, 7(1):??, January 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/01/o1003.pdf.

Goth:2007:CRB

[Got07a] Greg Goth. Chip researchers on brink of next-gen breakthroughs. IEEE Distributed Systems Online, 8(3):??, March 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/03/o3003.pdf.

Goth:2007:DLT

[Got07b] Greg Goth. Digital libraries are taking form. IEEE Distributed Systems Online, 8(12): ??, December 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/12/
                                        mds2007120004.pdf.

Goth:2007:FTS

[Got07c] Greg Goth. Faster than smoke? more blinding than mirrors? IEEE Distributed Systems Online, 8(7):??, July 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/07/
mds2007070005.pdf.

Goth:2007:KMS

[Got07d] Greg Goth. Key management standards hit the fast track. IEEE Distributed Systems Online, 8(9):??, September 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/09/mds2007090004.pdf.

Goth:2007:NSU

[Got07e] Greg Goth. News: a structure for unstructured data search. IEEE Distributed Systems Online, 8(1):??, January 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/01/o1003.pdf.

Goth:2007:OMN

[Got07f] Greg Goth. Opening the mobile net. IEEE Distributed Systems Online, 8(11):??, November 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/11/mds2007110004.pdf.

Goth:2007:PDA

[Got07g] Greg Goth. The politics of DDoS attacks. IEEE Distributed Systems Online, 8 (8):??, August 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/08/
                                        mds2007080003.pdf.

Goth:2007:SRM

[Got07h] Greg Goth. “ring-ring” might soon mean ka-ching. IEEE Distributed Systems Online, 8(4): ??, April 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2007/04/o4004.pdf.

Goth:2007:RTM

[Got07i] Greg Goth. Royalties and taxes: Making money on the net. IEEE Distributed Systems Online, 8(6):??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/06/o6006.pdf.

Goth:2007:SLT

[Got07j] Greg Goth. Stepping up to long-term research. IEEE Distributed Systems Online, 8 (10):??, October 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/10/
                                        mds2007100003.pdf.

Goth:2007:VOT

[Got07k] Greg Goth. Virtualization: Old technology offers huge new potential. IEEE Distributed Systems Online, 8(2):??, February 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/02/o2003.pdf.

Goth:2007:YH

[Got07l] Greg Goth. You are here. IEEE Distributed Systems Online, 8 (5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/05/o5007.pdf.

Goth:2008:FCR

[Got08a] G. Goth. Financial crisis and the role of risk-management software. IEEE Distributed Systems Online, 9(11):??, November 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:ITM

[Got08b] G. Goth. ISP traffic management: Will innovation or regulation ensure fairness? IEEE Distributed Systems Online, 9 (9):??, September 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:NWF

[Got08c] G. Goth. New wi-fi technology racing past standards process. IEEE Distributed Systems Online, 9(10):??, October 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:SSSb

[Got08d] G. Goth. Single sign-on and social networks. IEEE Distributed Systems Online, 9(12):??, December 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:SSSa

[Got08e] G. Goth. Software-as-a-service: The spark that will change software engineering? IEEE Distributed Systems Online, 9(7): ??, July 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:WWB

[Got08f] G. Goth. When Web 2.0 becomes Web Uh-Oh. IEEE Distributed Systems Online, 9(8): ??, August 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Goth:2008:SNS

[Got08g] Greg Goth. Are social networking sites growing up? IEEE Distributed Systems Online, 9 (2):??, February 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/02/
mds2008020003.pdf.

Goth:2008:BCS

[Got08h] Greg Goth. Biometrics could streamline border crossings. IEEE Distributed Systems Online, 9(4):??, April 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/04/
mds2008040002.pdf.

Goth:2008:N

[Got08i] Greg Goth. News. IEEE Distributed Systems Online, 9 (1):??, January 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/01/
mds2008010002.pdf.

Goth:2008:PP

[Got08j] Greg Goth. Power to the patients. IEEE Distributed Systems Online, 9(5):??, May 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/05/mds2008050001.pdf.

Goth:2008:TIC

[Got08k] Greg Goth. Terror on the Internet: a complex issue, and getting harder. IEEE Distributed Systems Online, 9(3):??, March 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/03/mds2008030003.pdf.

Goth:2008:ULC

[Got08l] Greg Goth. Ultra low-cost PCs redraw the OS wars. IEEE Distributed Systems Online, 9(6):??, June 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/06/
mds2008060003.pdf.

Guessoum:2004:AAM

[Gue04] Z. Guessoum. Adaptive agents and multiagent systems. IEEE Distributed Systems Online, 5 (7):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29290/01323039.pdf?isnumber=
29290&prod=JNL&arnumber=1323039&
arSt=&ared=&arAuthor=Guessoum%
2C+Z.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29290&arnumber=1323039&
count=4&index=1;   http:/
/ieeexplore.ieee.org/xpls/
references.jsp?arnumber=1323039.

Hill:2004:MAS

[HAMC^{+}04] R. Hill, J. Al-Muhtadi, R. Campbell, A. Kapadia, P. Naldurg, and A. Ranganathan. A middleware architecture for securing ubiquitous computing cyber infrastructures. IEEE Distributed Systems Online, 5(9):1, September 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29771/01355901.pdf?isnumber=
29771&prod=JNL&arnumber=1355901&
arSt=+1&ared=+1&arAuthor=+
Hill%2C+R.%3B++Al-Muhtadi%
2C+J.%3B++Campbell%2C+R.%
3B++Kapadia%2C+A.%3B++Naldurg%
2C+P.%3B++Ranganathan%2C+
A.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29771&arnumber=1355901&count=
4&index=0.

Hughes:2005:FRG

[HCW05] Daniel Hughes, Geoff Coulson, and James Walkerdine. Free riding on Gnutella revisited: The bell tolls? IEEE Distributed Systems Online, 6(6):??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2005/06/o6001.pdf.

Hedberg:2000:BDC

[Hed00a] Sara Hedberg. Beyond desktop computing: MIT’s Oxygen Project. IEEE Distributed Systems Online, 1 (1), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds100/ds1newprint.
htm.

Hedberg:2000:SIM

[Hed00b] Sara Reese Hedberg. Seattle’s Interactive Music Museum. IEEE Distributed Systems Online, 1(3), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds300/ds3newprint.
htm.

Hedberg:2001:ADC

[Hed01a]  Sara Reese Hedberg.  Af-

ter desktop computing: a progress report on smart environments research. IEEE Distributed Systems Online, 2 (1), 2001. CODEN ????

ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0101/highlights/spo0101_
print.htm.

Hedberg:2001:NGD

[Hed01b] Sara Reese Hedberg. A new generation of distance learning is emerging. IEEE Distributed Systems Online, 2 (2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0102/features/dis0102_
print.htm.

Halvorsen:2004:SSSb

[HGG^{+}04a] P. Halvorsen, C. Griwodz, V. Goebel, K. Lund, T. Plagemann, and J. Walpole. Storage system support for continuous media applications. 2. Multiple disks, memory, and integration. IEEE Distributed Systems Online, 5(2):1/1–1/16, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28454/01270781.pdf?isnumber=
28454&prod=JNL&arnumber=1270781&
arSt=+1%2F1&ared=+1%2F16&
arAuthor=+Halvorsen%2C+P.%
3B++Griwodz%2C+C.%3B++Goebel%
2C+V.%3B++Lund%2C+K.%3B++
Plagemann%2C+T.%3B++Walpole%
2C+J.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28454&arnumber=1270781&
count=6&index=0.

Halvorsen:2004:SSSa

[HGG^{+}04b] P. Halvorsen, C. Griwodz, V. Goebel, K. Lund, T. Plagemann, and J. Walpole. Storage system support for continuousmedia applications. Part 1. Requirements and single-disk issues. IEEE Distributed Systems Online, 5(1):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28452/01270711.pdf?isnumber=
28452&prod=JNL&arnumber=1270711&
arSt=&ared=&arAuthor=Halvorsen%
2C+P.%3B+Griwodz%2C+C.%3B+
Goebel%2C+V.%3B+Lund%2C+K.
%3B+Plagemann%2C+T.%3B+Walpole%
2C+J.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28452&arnumber=1270711&
count=8&index=1;   http:/
/ieeexplore.ieee.org/xpls/
references.jsp?arnumber=1270711.

Heflin:2003:IIC

[HH03] Jeff Heflin and Michael N.

Huhns. IEEE Internet computing: Guest Editors’ introduction: The zen of the Web. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/f/w5geip.htm.

Hilf:2003:BRP

[Hil03]  Bill Hilf.  Book review:

The persuasive power of computers A review of Persuasive Technology: Using Computers to Change What we Think and Do by B. J. Fogg. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0306/d/bks_a.htm.

Hedin:2007:CPM

[HL07] Björn Hedin and Erik Lindgren. A comparison of presentation methods for reading on mobile phones. IEEE Distributed Systems Online, 8(6):??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/06/o6002.pdf.

Hadim:2006:MMC

[HM06] Salem Hadim and Nader Mohamed. Middleware: Middleware challenges and approaches for wireless sensor networks. IEEE Distributed Systems Online, 7(3):??, March 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/03/
                                        o3001.pdf.

Herrmann:2005:SMS

[HMG05] K. Herrmann, G. Muhl, and K. Geihs. Self management: the solution to complexity or just another problem? IEEE Distributed Systems Online, 6(1): ??, January 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/30521/01407752.pdf?isnumber=
30521&prod=JNL&arnumber=1407752&
arSt=&ared=&arAuthor=Herrmann%
2C+K.%3B+Muhl%2C+G.%3B+Geihs%
2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30521&arnumber=1407752&
count=4&index=0.

Horre:2008:DAM

[HMJV08] Wouter Horré, Sam Michiels, Wouter Joosen, and Pierre Verbaeten. DAVIM: Adaptable middleware for sensor networks. IEEE Distributed Systems Online, 9(1):??, January 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/01/mds2008010001.pdf.

Huhns:2003:IIC

[Huh03] Michael N. Huhns. IEEE Internet computing: Agents on the Web: The Sentient Web. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/d/w6age.htm.

Hughes:2006:PPD

[HWCG06] Daniel Hughes, James Walkerdine, Geoff Coulson, and Stephen Gibson. Peer-to-peer: Is deviant behavior the norm on P2P file-sharing networks? IEEE Distributed Systems Online, 7(2):??, February 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/02/o2001.pdf.

Ionescu:2001:LCR

[IM01] Mihail Ionescu and Ivan Marsic. Latecomer and crash recovery support in fault-tolerant groupware. IEEE Distributed Systems Online, 2(7), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0107/features/ionescu_
                                        final.pdf.

Ivanovic:2003:BRP

[Iva03] Vladimir G. Ivanovíc. Book review: The power of events: a review of The Power of Events: An Introduction to Complex Event Processing in Distributed Enterprise Systems by David Luckham. IEEE Distributed Systems Online, 4(3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0303/d/bks_d.htm.

Jain:2002:FC

[Jai02] Ramesh Jain. Folk computing. IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0208/d/med.htm.

Jain:2007:TE

[Jai07] Ramesh Jain. Toward EventWeb. IEEE Distributed Systems Online, 8(9):??, September 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/09/mds2007090002.
pdf.

Jakubczyc:2002:MSS

[Jak02] Janina Jakubczyc. Motivating self-stabilization: a review of Self-Stabilization by Shlomi Dolev. IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/d/bks_b.htm.

Jaeger:2008:FBI

[JG08] M. C. Jaeger and M. Gero. A floating backbone for Internet over the ocean. IEEE Distributed Systems Online, 9(12): ??, December 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Joita:2006:ADC

[JRC^{+}06] Liviu Joita, Omer F. Rana, Pablo Chacín, Isaac Chao, Felix Freitag, Leandro Navarro, and Oscar Ardaiz. Application deployment on catallactic Grid middleware. IEEE Distributed Systems Online, 7(12):??, December 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/12/oz001.pdf.

Kaczmarek:2007:BR

[Kac07] Tomasz Kaczmarek. Book reviews. IEEE Distributed Systems Online, 8(12):??, December 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/12/mds2007120005.
pdf.

Karatza:2002:BRO

[Kar02] Helen Karatza. Book review: An owner’s manual for high performance computing: a review of Techniques for Optimizing Applications: High Performance Computing by Rajat P. Garg and Ilya Sharapov. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/bks_a.htm.

Karatza:2003:BRP

[Kar03]  Helen Karatza.  Book re-

view: pH: a language for implicit parallel programming: a review of Implicit Parallel Programming in pH by Rishiyur S. Nikhil & Arvind. IEEE Distributed Systems Online, 4(4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/d/bks_b.htm.

Karatza:2004:ERP

[Kar04a]  H. Karatza.  An excellent

resource for parallel computing. IEEE Distributed Systems Online, 5(7):5, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/29290/01323040.pdf?isnumber=
29290&prod=JNL&arnumber=1323040&
arSt=+5&ared=+5&arAuthor=
Karatza%2C+H.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29290&arnumber=1323040&
count=4&index=2.

Karatza:2004:VRP

[Kar04b] H. Karatza. A valuable resource of programmers using .NET technology. IEEE Distributed Systems Online, 5(6): 4, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29028/01308176.pdf?isnumber=
29028&prod=JNL&arnumber=1308176&
arSt=+4&ared=+4&arAuthor=
Karatza%2C+H.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29028&arnumber=1308176&
count=2&index=1.

Karatza:2004:ISA

[Kar04c] H. D. Karatza. Intelligent software agents’ influence on mediabased systems. IEEE Distributed Systems Online, 5(3): 1–3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/

                                        8968/28687/01285881.pdf?isnumber=
                                        28687&prod=JNL&arnumber=1285881&
                                        arSt=+1&ared=+3&arAuthor=
                                        Karatza%2C+H.D.;   http:/
                                        /ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=28687&
                                        arnumber=1285881&count=7&
                                        index=4.

Kapitza:2006:FIA

[KDH^{+}06] Rüdiger Kapitza, Jörg Domaschka, Franz J. Hauck, Hans P. Reiser, and Holger Schmidt. FORMI: Integrating adaptive fragmented objects into Java RMI. IEEE Distributed Systems Online, 7(10): ??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/10/ox001.pdf.

Kotz:2002:FDM

[KGR02] David Kotz, Robert Gray, and Daniela Rus. Future directions for mobile agent research. IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0208/f/kot_print.htm.

Kilmer:2002:PBP

[Kil02] Richard Kilmer. Peering beyond the PC: Where P2P meets the wireless Web. IEEE Distributed Systems Online, 3 (5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.

org/0205/features/w3icon.
htm.

Kim:2002:FTD

[Kim02] K. H. (Kane) Kim. Faulttolerant distributed computing: Evolution and issues. IEEE Distributed Systems Online, 3(7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/features/kim_print.
htm.

Kim:2003:BRB

[Kim03]  Maksim Kim.  Book re-

view: Building Web-integrated database applications: a review of Web Database Applications with PHP & MySQL by Hugh E. Williams and David Lane. IEEE Distributed Systems Online, 4(2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/d/bks_a.htm.

Kurdi:2008:CET

[KLAR08] Heba Kurdi, Maozhen Li, and Hamed Al-Raweshidy. A classification of emerging and traditional Grid systems. IEEE Distributed Systems Online, 9 (3):??, March 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/03/
mds2008030001.pdf.

Krishnaswamy:2004:ECT

[KLZ04] S. Krishnaswamy, S. W. Loke, and A. Zaslavsky. Estimating computation times of dataintensive applications. IEEE Distributed Systems Online, 5 (4):??, April 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/28913/01301253.
pdf?isnumber=28913&prod=JNL&
arnumber=1301253&arSt=&ared=
&arAuthor=Krishnaswamy%2C+
S.%3B+Loke%2C+S.W.%3B+Zaslavsky%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28913&arnumber=1301253&
count=5&index=0.

Kordon:2007:GEIa

[Kor07a] Fabrice Kordon. Guest Editor’s introduction: Rapid system prototyping. IEEE Distributed Systems Online, 8(3):??, March 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/03/o3007.pdf.

Kordon:2007:GEIb

[Kor07b] Fabrice Kordon. Guest Editor’s introduction: Rapid system prototyping. IEEE Distributed Systems Online, 8(4):??, April 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/04/o4007.pdf.

Kordon:2007:GEIc

[Kor07c] Fabrice Kordon. Guest Editor’s introduction: Rapid system prototyping. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/05/o5001.pdf.

Kornecki:2008:CCC

[Kor08] Andrew J. Kornecki. Computing curricula for the 21st Century. IEEE Distributed Systems Online, 9(2):??, February 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/02/mds2008020002.pdf.

Kowalkiewicz:2006:BRU

[Kow06] Marek Kowalkiewicz. Book review: a useful but inadequate introduction to Web services. IEEE Distributed Systems Online, 7(6):??, June 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/06/o6006.pdf.

Kordon:2005:TNG

[KP05] F. Kordon and L. Pautet. Toward nex-generation middleware? IEEE Distributed Systems Online, 6(3):??, March 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30740/01423917.pdf?isnumber=
30740&prod=JNL&arnumber=1423917&
arSt=&ared=&arAuthor=Kordon%
2C+F.%3B+Pautet%2C+L.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30740&
arnumber=1423917&count=4&
index=1.

Kordon:2006:TFM

[KP06] Fabrice Kordon and Laure Petrucci. Toward formalmethods oecumenism? IEEE Distributed Systems Online, 7 (7):??, July 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/07/o7002.pdf.

Khan:2008:ACM

[KRG08] M. U. Khan, R. Reichle, and K. Geihs. Architectural constraints in the model-driven development of self-adaptive applications. IEEE Distributed Systems Online, 9(7):??, July 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/xpls/
abs_all.jsp?arnumber=4589947.

Kulik:2007:MCS

[Kul07] Lars Kulik. Mobile computing systems programming: a graduate distributed computing course. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.

computer.org/comp/mags/ds/
2007/05/o5004.pdf.

Kumar:2003:EPC

[Kum03] Mohan Kumar. Education:

Pervasive computing education. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/d/d7edu_print.htm.

Lathia:2005:AGC

[Lat05a] M. Lathia. Advantages of Grid computing. IEEE Distributed Systems Online, 6(2): 5, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/30522/01407762.pdf?isnumber=
30522&prod=JNL&arnumber=1407762&
arSt=+5&ared=+5&arAuthor=
Lathia%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30522&arnumber=1407762&
count=5&index=3.

Lathia:2005:URP

[Lat05b] M. Lathia. A useful resource for parallel and distributed computing. IEEE Distributed Systems Online, 6(4):4, April 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30778/01425407.pdf?isnumber=
30778&prod=JNL&arnumber=1425407&
arSt=+4&ared=+4&arAuthor=+
Lathia%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?

                                        isnumber=30778&arnumber=1425407&
                                        count=3&index=1.

Lathia:2006:BRG

[Lat06a] Milan Lathia. Book review: Grid technology concepts from A to Z. IEEE Distributed Systems Online, 7(12):??, December 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/12/oz004.pdf.

Lathia:2006:BRN

[Lat06b] Milan Lathia. Book reviews: The new wave — concurrent programming. IEEE Distributed Systems Online, 7(4): ??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/04/o4005.pdf.

Lathia:2007:CBC

[Lat07] Milan Lathia. Computing in bioinformatics and computational biology: a collection. IEEE Distributed Systems Online, 8(3):??, March 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/03/
                                        o3004.pdf.

Lawson:2004:VSW

[Law04]  K. Lawson.  Visualizing

the Semantic Web: XMLbased Internet and Information Visualization [book review]. IEEE Distributed Systems Online, 5(2):6/1–6/3, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28454/01270786.pdf?isnumber=
28454&prod=JNL&arnumber=1270786&
arSt=+6%2F1&ared=+6%2F3&arAuthor=
+Lawson%2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28454&arnumber=1270786&
count=6&index=4.

Low:2008:DCE

[LCF^{+}08] B. Low, K. Cassidy, D. Fergusson, M. Atkinson, E. V. Meer, and M. McGeever. Distributed computing education, part 5: Coming to terms with intellectual property rights. IEEE Distributed Systems Online, 9(12): ??, December 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Lewin:2002:SB

[Lew02a]  David Lewin.  Strike up

the broadband. IEEE Distributed Systems Online, 3 (2), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0202/departments/new_
print.htm.

Lewin:2002:SDS

[Lew02b]  David I. Lewin.  Six de-

grees of searches. IEEE Distributed Systems Online, 3 (6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0206/features/new_1.htm.

Lee:2003:MDM

[LHWX03] Dik Lun Lee, Ka Kit Hoi, Wing Sing Dicky Wong, and Jianliang Xu. Mobile data management: Web browsing on small displays. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/f/leep.htm.

Liebrock:2006:ESS

[Lie06] Lorie M. Liebrock. Education: Scholarship for service. IEEE Distributed Systems Online, 7(9):??, September 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/09/o9002.pdf.

Liebrock:2007:SSN

[Lie07] Lorie M. Liebrock. Scholarship for service: New Mexico Tech’s undergraduate program. IEEE Distributed Systems Online, 8(3):??, March 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/03/o3002.pdf.

Lingaiah:2002:BRC

[Lin02] Dharmendra Lingaiah. Book review: Communication sans wires: a review of Wireless Communications and Networks by William Stallings. IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/d/bks_a.htm.

Lingaiah:2003:BRD

[Lin03a] Dharmendra Lingaiah. Book review: Data dependency and optimization: a review of Optimizing Compilers for Modern Architecture by Randy Allen and Ken Kennedy. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0306/d/bks_b.htm.

Lingaiah:2003:BRO

[Lin03b] Dharmendra Lingaiah. Book review: OFDM wireless communications: a review of OFDM for Wireless Multimedia Communications by Richard Van Nee and Ramjee Prasad. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/0308/d/bks_c.htm.

Lingaiah:2003:BRT

[Lin03c] Dharmendra Lingaiah. Book review: Telecom foundations: a review of telecommunications essentials by Lillian Goleniewski. IEEE Distributed Systems Online, 4 (1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0301/d/bks_a.htm.

Liu:2006:BRS

[Liu06] Nelson Liu. Book reviews: a summary of multifractal traffic modeling. IEEE Distributed Systems Online, 7(3):??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/03/o3008.pdf.

Liu:2007:MWP

[Liu07] Ed Curry Knarig Arabshian Henning Schulzrinne Mohammed Jubaer Arif Shanika Karunasekera Santosh Kulkarni Roberto Speicys Cardoso Sonia Ben-Mokhtar Valérie Issarny Aitor Urbieta Manos Kapritsos Peter Triantafillou Juan López Rubio Hao Yang Fan Ye Zhen Liu. Middleware 2007 works in progress. IEEE Distributed Systems Online, 8(11): ??, November 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/11/
                                        mds2007110003.pdf.

Lutz:2003:ISR

[LL03] Michael H. Lutz and Phillip A. Laplante. IEEE Software: Real-time systems: C# and the .NET framework: Ready for real time? IEEE Distributed Systems Online, 4 (2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/f/sp1lap.htm.

Llorente:2005:LCV

[LMH05] I. M. Llorente, R. S. Montero, and E. Huedo. A loosely coupled vision for computational Grids. IEEE Distributed Systems Online, 6(5):2, May 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/31212/01453471.pdf?isnumber=
31212&prod=JNL&arnumber=1453471&
arSt=+2&ared=+2&arAuthor=+
Llorente%2C+I.M.%3B++Montero%
2C+R.S.%3B++Huedo%2C+E.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
31212&arnumber=1453471&count=
4&index=1.

Liljenstam:2001:MPS

[LRA01] Michael Liljenstam, Robert Rönngren, and Rassul Ayani. MobSim++: Parallel simulation of personal communication networks. IEEE Distributed Systems Online, 2 (2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0102/features/lil0102_
print.htm.

Linden:2003:IRA

[LSY03] Greg Linden, Brent Smith, and Jeremy York. Industry report: Amazon.com recommendations: Item-to-item collaborative filtering. IEEE Distributed Systems Online, 4 (1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0301/d/wp1lind.htm.

Lea:2003:IIC

[LV03] Doug Lea and Steve Vinoski. IEEE Internet computing: Guest Editors’ introduction: Middleware for Web services. IEEE Distributed Systems Online, 4(1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0301/f/w1gei.htm.

Manzanares:2005:SNP

[Man05] A. I. Manzanares. Sensible network protection. IEEE Distributed Systems Online, 6(5): 6, May 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/31212/01453475.pdf?isnumber=
                                        31212&prod=JNL&arnumber=1453475&
                                        arSt=+6&ared=+6&arAuthor=+
                                        Manzanares%2C+A.I.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=31212&
                                        arnumber=1453475&count=4&
                                        index=3.

Manzanares:2007:HNI

[Man07] Antonio Izquierdo Manzanares. Hardening network infrastructure: Not suitable for everyone. IEEE Distributed Systems Online, 8(10):??, October 2007.

CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/10/mds2007100004.pdf.

Marin:2005:WCS

[Mar05] O. Marin. So where’s the computation superhighway? [WWW computing resources]. IEEE Distributed Systems Online, 6(1):??, January 2005. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://ieeexplore.ieee.
org/iel5/8968/30521/01407753.
pdf?isnumber=30521&prod=JNL&
arnumber=1407753&arSt=&ared=
&arAuthor=Marin%2C+O.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30521&
arnumber=1407753&count=4&
index=1.

Marowka:2008:TPT

[Mar08a] A. Marowka. Think parallel:

Teaching parallel programming today. IEEE Distributed Systems Online, 9(8):??, August 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Marowka:2008:BR

[Mar08b] Ami Marowka. Book review.

IEEE Distributed Systems Online, 9(1):??, January 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/01/mds2008010003.pdf.

Mossiere:2006:WP

[MCB^{+}06] Jacques Mossière, Edward Curry, Abdelkrim Beloued, Jean-Marie Gilliot, Françoise André, Maria-Teresa Segarra, Vladimir Dyo, Etienne Antoniutti Di Muro, and Michael A. Jaeger. Works in progress. IEEE Distributed Systems Online, 7(2):??, February 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/02/o2003.pdf.

McGinity:2003:BBG

[McG03a] Meg McGinity. In brief: Bullies get technical. IEEE Distributed Systems Online, 4 (2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0302/d/brief.htm.

McGinity:2003:BLL

[McG03b] Meg McGinity. In brief: Live and e-learn? IEEE Distributed Systems Online, 4 (1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0301/d/brief1.htm.

McGinity:2003:NDS

[McG03c] Meg McGinity. News: Datastoring DNA. IEEE Distributed Systems Online, 4 (4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/f/news.htm.

McGinity:2003:NRY

[McG03d] Meg McGinity. News: RFID:

Not your father’s bar code. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/f/newsp.htm.

McLaughlin:2002:BLC

[McL02]  Laurianne McLaughlin.  In

brief: Library of Congress tests Grid technology. IEEE Distributed Systems Online, 3 (12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0212/d/brief2.htm.

McLaughlin:2003:NNA

[McL03a] Laurianne McLaughlin. News:

Net access through powerlines: The new buzz in broadband. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/f/news_pfv.htm.

McLaughlin:2003:NVDa

[McL03b] Laurianne McLaughlin. News:

Virginia draws a new line on spam. IEEE Distributed Systems Online, 4(6), 2003. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://dsonline.computer.
org/0306/f/news.htm.

McLaughlin:2003:NVDb

[McL03c] Laurianne McLaughlin. News: VoIP draws new scrutiny from FBI, State regulators. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/f/newsp.htm.

McLaughlin:2004:WNW

[McL04] L. McLaughlin. What’s next in Web search? IEEE Distributed Systems Online, 5(11): 2, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/30052/01377089.
pdf?isnumber=30052&prod=JNL&
arnumber=1377089&arSt=+2&
ared=+2&arAuthor=+McLaughlin%
2C+L.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30052&arnumber=1377089&
count=5&index=1.

McLaughlin:2006:BWW

[McL06] Laurianne McLaughlin. In brief: Where wired and wireless networks converge. IEEE Distributed Systems Online, 7(5): ??, May 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/05/o5005.pdf.

Mehra:2002:ICG

[Meh02] Pankaj Mehra. Internet computing: Guest Editor’s introduction: Global deployment of data centers. IEEE Distributed Systems Online, 3 (10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/f/wp5gei.htm.

Metz:2003:IIC

[Met03]  Chris Metz.  IEEE Inter-

net computing: Guest Editor’s introduction: Moving toward an IPv6 future. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/f/w3gei.htm.

McGuinness:2002:IIS

[MFHS02] Deborah L. McGuinness, Richard Fikes, James Hendler, and Lynn Andrea Stein. IEEE intelligent systems: DAML + OIL: An ontology language for the Semantic Web. IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/f/news_print.htm.

Midkiff:2002:NCS

[Mid02] Scott F. Midkiff. Networkcentric systems. IEEE Distributed Systems Online, 3 (6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0206/departments/bp2edu.
                                        htm.

Milojicic:2000:IT

[Mil00a] Dejan Milojicic. Internet technology. IEEE Distributed Systems Online, 1(3), 2000. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://dsonline.computer.
                                        org/archives/ds300/ds3treprint.
                                        htm.

Milojicic:2000:IFB

[Mil00b] Dejan Milojicic. Interview: Fred B. Schneider on distributed computing. IEEE Distributed Systems Online, 1 (1), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds100/ds1intprint.
                                        htm.

Milojicic:2002:DLL

[Mil02] Dejan Milojicic. A discussion with Leslie Lamport. IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0208/f/lam_print.htm;
                                         http://lamport.org/.

Milojicic:2004:MJG

[Mil04] D. Milojicic. Microsoft’s Jim Gray on computing’s breakthroughs, lessons, and future. IEEE Distributed Systems Online, 5(1):4.1–4.8, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28452/01270713.pdf?isnumber=
28452&prod=JNL&arnumber=1270713&
arSt=+4.1&ared=+4.8&arAuthor=
Milojicic%2C+D.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28452&
arnumber=1270713&count=8&
index=2.

Milojicic:2005:BLP

[Mil05] D. S. Milojicic. Barbara Liskov on programming, career, and the future. IEEE Distributed Systems Online, 6(2):2, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30522/01407759.pdf?isnumber=
30522&prod=JNL&arnumber=1407759&
arSt=+2&ared=+2&arAuthor=
Milojicic%2C+D.S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30522&
arnumber=1407759&count=5&
index=1.

Milojicic:2007:SMIa

[Mil07a] Dejan Milojicic. Service management: An interview with Alan Ganek. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/05/o5005.pdf.

Milojicic:2007:SMIb

[Mil07b] Dejan Milojicic. Service management: An interview with Mark Potts. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/05/o5006.pdf.

Milojicic:2007:WEI

[Mil07c] Dejan Milojicic. Web 2.0 and Enterprise IT. IEEE Distributed Systems Online, 8(12): ??, December 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/12/
                                        mds2007120003.pdf.

Mirhosseini:2002:CCA

[Mir02] Ali R. Mirhosseini. Computers as components: The art of embedded systems programming: a review of Computers as Components: Principles of Embedded Computing System Design by Wayne Wolf. IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0208/d/bks_b.htm.

Mjolsnes:2002:MCR

[Mjø02] Stig F. Mjølsnes. Mobile commerce: a review of Mobile Commerce: Opportunities, Applications, and Technologies.

IEEE Distributed Systems Online, 3(4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/departments/bks_
b.htm.

Ming:2007:BGS

[MM07] Anlong Ming and Huadong Ma.

The biometrics Grid: a solution to biometric technologies. IEEE Distributed Systems Online, 8(9):??, September 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/09/mds2007090001.pdf.

Moore:2001:SCS

[Moo01] Arden Moore. Speeding computer software. IEEE Distributed Systems Online, 2 (4), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0104/features/new0104.
htm.

Morgan:2003:BRA

[Mor03] Donald Morgan. Book review:

Ant book full of practical advice: a review of Java Development with Ant by Eric Hatcher and Steve Loughran. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/d/bks_b.htm.

McKinley:2002:EID

[MSC^{+}02] Philip K. McKinley, Kurt Stirewalt, Betty H. C. Cheng, Laura K. Dillon, and Sandeep S. Kulkarni. Education: Interactive distributed applications and the computer science curriculum. IEEE Distributed Systems Online, 3 (10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/edu_print.htm.

Martins:2005:NPM

[MSCV05] P. Martins, P. Sousa, A. Casimiro, and P. Verissimo. A new programming model for dependable adaptive real-time applications. IEEE Distributed Systems Online, 6(5):1, May 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/31212/01453470.pdf?isnumber=
31212&prod=JNL&arnumber=1453470&
arSt=+1&ared=+1&arAuthor=
+Martins%2C+P.%3B++Sousa%
2C+P.%3B++Casimiro%2C+A.%
3B++Verissimo%2C+P.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=31212&
arnumber=1453470&count=4&
index=0.

Maier:2002:TTA

[MSPB02] Reinhard Maier, Georg Stöger, Stefan Poledna, and Günther Bauer. Time-triggered architecture: a consistent computing platform. IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/f/mai_print.htm.

Murray:2001:GTT

[Mur01] Shani Murray. Grid technology takes on the 21st Century. IEEE Distributed Systems Online, 2(2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0102/features/gri0102.
htm.

Murphy:2006:EHP

[Mur06] Tom Murphy. Education: Highperformance computing in community colleges? IEEE Distributed Systems Online, 7(4): ??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/04/o4003.pdf.

Murphy:2007:HPC

[Mur07] Tom Murphy. High-performance computing in high schools? IEEE Distributed Systems Online, 8(8):??, August 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/08/
mds2007080002.pdf.

Myerson:2001:RHG

[Mye01]  Judith M. Myerson.  Re-

view: a hands-on guide to TCP/IP networking. IEEE Distributed Systems Online, 2 (4), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0104/departments/bks0104b.
htm.

Naik:2003:DSW

[NABG03] Vinayak Naik, Anish Arora, Sandip Bapat, and Mohamed Gouda. Dependable systems: Whisper: Local secret maintenance in sensor networks. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/f/gaep.htm.

Nack:2003:IMM

[Nac03] Frank Nack. IEEE Multimedia: Media impact: Migrating from mobile telephony to multipurpose gadgets. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/d/up2med.htm.

Necat:2006:BRE

[Nec06] Beran Necat. Book reviews: An expert discussion of wireless technologies. IEEE Distributed Systems Online, 7(9): ??, September 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.

org/comp/mags/ds/2006/09/
o9005.pdf.

Necat:2007:WNH

[Nec07] Beran Necat. Wireless networking for home users. IEEE Distributed Systems Online, 8(5): ??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2007/05/o5008.pdf.

Neuwirth:2002:UW

[NR02] Christine M. Neuwirth and Susan Harkness Regli. Usability and the Web. IEEE Distributed Systems Online, 3 (4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/features/w2gei.htm.

Omma:2004:BMA

[Omm04]  M. Omma.  On build-

ing multithreaded applications. IEEE Distributed Systems Online, 5(4):1–3, April 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28913/01301256.pdf?isnumber=
28913&prod=JNL&arnumber=1301256&
arSt=+1&ared=+3&arAuthor=
Omma%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28913&arnumber=1301256&
count=5&index=3.

Othman:2001:DAC

[OOS01a] Ossama Othman, Carlos O’Ryan, and Douglas C. Schmidt. Designing an adaptive CORBA load balancing service using TAO. IEEE Distributed Systems Online, 2(4), 2001. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://dsonline.computer.
org/0104/features/oth0104_
print.htm.

Othman:2001:SCM

[OOS01b] Ossama Othman, Carlos O’Ryan, and Douglas C. Schmidt. Strategies for CORBA middlewarebased load balancing. IEEE Distributed Systems Online, 2 (3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0103/features/oth0103_
print.htm.

Oppliger:2003:IPM

[Opp03] Rolf Oppliger. IT professional: Microsoft Outlook Web access: Blessing or bane to security? IEEE Distributed Systems Online, 4(4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/f/fp1opp.htm.

Oberle:2006:WSS

[OSE06] Daniel Oberle, Steffen Staab, and Andreas Eberhart. Web systems: Semantic management of distributed Web applications. IEEE Distributed Systems Online, 7(5):??, May 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/05/o5001.pdf.

Ortiz-Tapia:2006:BRA

[OT06a] Arturo Ortiz-Tapia. Book review: An authentic breakthrough. IEEE Distributed Systems Online, 7(12):??, December 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/12/oz003.pdf.

Ortiz-Tapia:2006:BRE

[OT06b] Arturo Ortiz-Tapia. Book reviews: An essential point of departure. IEEE Distributed Systems Online, 7(10):??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/10/ox005.pdf.

Ortiz-Tapia:2006:PCF

[OT06c] Arturo Ortiz-Tapia. A primer on computational fields. IEEE Distributed Systems Online, 7 (8):??, August 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/08/
o8003.pdf.

Ortiz-Tapia:2007:BRD

[OT07a] Arturo Ortiz-Tapia. Book review: Data plus connectivity. IEEE Distributed Systems Online, 8(1):??, January 2007.

CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/01/o1004.pdf.

Ortiz-Tapia:2007:BRE

[OT07b] Arturo Ortiz-Tapia. Book reviews: An excellent Internet tutorial. IEEE Distributed Systems Online, 8(2):??, February 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/02/o2005.pdf.

Ortiz-Tapia:2007:DSA

[OT07c] Arturo Ortiz-Tapia. A door to state-of-the-art high-performance computing. IEEE Distributed Systems Online, 8(6):??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/06/o6004.pdf.

Ortiz-Tapia:2007:FIV

[OT07d] Arturo Ortiz-Tapia. A friendly introduction to VoIP. IEEE Distributed Systems Online, 8 (9):??, September 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/09/
mds2007090005.pdf.

Ortiz-Tapia:2007:RRL

[OT07e] Arturo Ortiz-Tapia. A recommended resource for Linuxbased companies. IEEE Distributed Systems Online, 8 (8):??, August 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/08/
mds2007080004.pdf.

Paprzycki:2006:EIP

[Pap06] Marcin Paprzycki. Education:

Integrating parallel and distributed computing in computer science curricula. IEEE Distributed Systems Online, 7 (2):??, February 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/02/
o2006.pdf.

Pardi:2004:PCD

[Par04] W. Pardi Jr. Programming concurrent and distributed algorithms in Java. IEEE Distributed Systems Online, 5(11): 5, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/30052/01377092.
pdf?isnumber=30052&prod=JNL&
arnumber=1377092&arSt=+5&
ared=+5&arAuthor=+Pardi%2C+
W.%2C+Jr.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30052&arnumber=1377092&
count=5&index=4.

Pawlak:2006:SCT

[Paw06]  Renaud Pawlak.  Spoon:

Compile-time annotation processing for middleware. IEEE Distributed Systems Online, 7 (11):??, November 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/11/
                                        oy001.pdf.

Piva:2002:MCW

[PBB02] Alessandro Piva, Franco Bartolini, and Mauro Barni. Managing copyright: Watermark and cryptography algorithms. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0205/features/w3piva.
                                        htm.

Perez:2003:BRC

[Pér03a] Hugo Jiménez Pérez. Book review: a concise and simple approach for LDAP beginners: a review of LDAP Programming, Management, and Integration by Clayton Donley. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0309/d/bks_b.htm.

Perez:2003:BRN

[Pér03b] Mat Hugo Jiménez Pérez. Book review: Network+ certification: a good starting point: a review of Network+ Certification Training Kit by Microsoft. IEEE Distributed Systems Online, 4(1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0301/d/bks_c.htm.

Petcu:2005:PNA

[Pet05] D. Petcu. Parallel numerical applications on Grid environments. IEEE Distributed Systems Online, 6(4):5, April 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30778/01425408.pdf?isnumber=
30778&prod=JNL&arnumber=1425408&
arSt=+5&ared=+5&arAuthor=+
Petcu%2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30778&arnumber=1425408&
count=3&index=2.

Petcu:2008:CDG

[Pet08a] Dana Petcu. A comprehensive development guide for the Globus Toolkit. IEEE Distributed Systems Online, 9(6):??, June 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2008/06/
mds2008060004.pdf.

Petcu:2008:TGTa

[Pet08b] Dana Petcu. Teaching Grid technologies to PhD students, part 1: Using best practices to build the course. IEEE Distributed Systems Online, 9 (3):??, March 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2008/03/
                                        mds2008030002.pdf.

Petcu:2008:TGTb

[Pet08c] Dana Petcu. Teaching Grid technologies to PhD students, part 2: Course structure and experiences. IEEE Distributed Systems Online, 9(4):??, April 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2008/04/mds2008040001.pdf.

Pham:2001:PWL

[Pha01] Hanh Pham. A practical way to learn modern applications design. IEEE Distributed Systems Online, 2(2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0102/departments/bks0102.
                                        htm.

Polk:2003:IIC

[PHM03] William T. Polk, Nelson E. Hastings, and Ambarish Malpani. IEEE Internet computing: Security track: Public key infrastructures that satisfy security goals. IEEE Distributed Systems Online, 4 (7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0307/f/wp4sec.htm.

Pierre:2006:EWS

[Pie06] Guillaume Pierre. From the Editor: Web systems community — to those who believe we are done already. IEEE Distributed Systems Online, 7(1): ??, January 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/01/o1002.pdf.

Parzyjegla:2008:MDD

[PJMW08] H. Parzyjegla, M. A. Jaeger, G. Muhl, and T. Weis. Modeldriven development and adaptation of autonomous control applications. IEEE Distributed Systems Online, 9(11): ??, November 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Pashtan:2003:IIC

[PKP03] Ariel Pashtan, Shriram Kollipara, and Michael Pearce. IEEE Internet computing: Industry report: Adapting content for wireless Web services. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/d/w5indp.htm.

Pedro:2007:SPV

[PLB07] Luis Pedro, Levi Lucio, and Didier Buchs. System prototype and verification using metamodel-based transformations. IEEE Distributed Systems Online, 8(4):??, April 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.

                                        computer.org/comp/mags/ds/
                                        2007/04/o4001.pdf.

Pietzuch:2007:DEB

[PMF07] Peter Pietzuch, Gero Mühl, and Ludger Fiege. Distributed event-based systems: An emerging community. IEEE Distributed Systems Online, 8 (2):??, February 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/02/
                                        o2002.pdf.

Patterson:2003:IPCa

[PMP03a] Cynthia A. Patterson, Richard R. Muntz, and Cherri M. Pancake. IEEE pervasive computing: Spotlight: Challenges in location-aware computing. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0306/d/b2spo.htm.

Patterson:2003:IPCb

[PMP03b] Cynthia A. Patterson, Richard R. Muntz, and Cherri M. Pancake. IEEE pervasive computing: Spotlight: Challenges in location-aware computing. IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0307/d/bp2spo.htm.

Patterson:2003:IPCc

[PMP03c] Cynthia A. Patterson, Richard R.

Muntz, and Cherri M. Pancake. IEEE pervasive computing: Spotlight: Challenges in location-aware computing. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/d/bp2spo.htm.

Patterson:2003:IPCd

[PMP03d] Cynthia A. Patterson, Richard R.

Muntz, and Cherri M. Pancake. IEEE pervasive computing: SPOTLIGHT challenges in location-aware computing. IEEE Distributed Systems Online, 4(9), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0309/d/b2spo.htm.

Patterson:2003:IPCe

[PMP03e] Cynthia A. Patterson, Richard R.

Muntz, and Cherri M. Pancake. IEEE pervasive computing: Spotlight: Challenges in location-aware computing. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0310/d/bspo.htm.

Patil:2003:IIC

[PN03] Sanjay Patil and Eric Newcomer. IEEE Internet computing: Spotlight ebXML and Web services. IEEE Distributed Systems Online, 4 (5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/f/wp3spot.htm.

Powers:2001:WPP

[PO01] Paula Powers and Vladimir Oboronko. The Wired for Peace Project. IEEE Distributed Systems Online, 2 (2), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0102/features/wir0102_
                                        print.htm.

Powers:2001:ICT

[Pow01] Paula Powers. Internet connections 1,000 times faster: DARPA’s SuperNet. IEEE Distributed Systems Online, 2 (1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0101/features/dar0101.
                                        htm.

Prakash:2001:MC

[Pra01] Ravi Prakash. Mobile computing. IEEE Distributed Systems Online, 2(6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0106/departments/edu0106_
                                        print.htm.

Pankanti:2003:IPC

[PSB^{+}03] Sharath Pankanti, Andrew Senior, Lisa Brown, Arun Hampapur, Ying-Li Tian, Ruud Bolle, Ma Celeste Campo Florina Almeniárez, Andrés Marín, Carlos García R., and Rob van Kranenburg. IEEE pervasive computing: Works in progress: Security, privacy, and health: PeopleVision: Privacy protection in visual surveillance; security model for intelligent-agents-based pervasive computing environments; ubicomp to provide feedback for people with obsessive-compulsive disorder. IEEE Distributed Systems Online, 4(3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0303/f/b1wip.htm.

Payne:2002:CAS

[PSS02] Terry R. Payne, Rahul Singh, and Katia Sycara. Calendar agents on the semantic Web. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0205/departments/sem.
htm.

Puzmanova:2002:BRW

[Puz02a] Rita Puzmanova. Book review: Wireless Web: Not just for managers: a review of Wireless Web: A Manager’s Guide by Frank P. Coyle.

IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0211/d/bks_c.htm.

Puzmanova:2002:IAW

[Puz02b] Rita Puzmanova. The Internet for all who are curious: a review of The Internet Book: Everything You Need to Know About Computer Networking and How the Internet Works, 3rd ed., by Douglas E. Comer. IEEE Distributed Systems Online, 3(9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0209/d/bks_a.htm.

Puzmanova:2003:BRD

[Puz03a] Rita Puzmanova. Book review: Delivering voice over IP and multimedia with session initiation protocol: a review of Internet Communications Using SIP by Henry Sinnreich and Alan B. Johnston. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/d/bks_a.htm.

Puzmanova:2003:BRI

[Puz03b] Rita Puzmanova. Book review: Internet for the mobile world: a review of The Wireless Mobile Internet: Architectures, Protocols and Services by Abbas Jamalipour. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/d/bks_b.htm.

Puzmanova:2003:BRL

[Puz03c] Rita Puzmanova. Book review:

Learning about WDM: a review of IP Over WDM by Kevin H. Liu. IEEE Distributed Systems Online, 4(4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/d/bks_a.htm.

Puzmanova:2004:HNB

[Puz04] R. Puzmanova. Home Networking Basis: Transmission Environments and Wired/ Wireless Protocols [book review]. IEEE Distributed Systems Online, 5(2):7/1–7/3, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28454/01270787.pdf?isnumber=
28454&prod=JNL&arnumber=1270787&
arSt=+7%2F1&ared=+7%2F3&arAuthor=
+Puzmanova%2C+R.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270787&count=6&
index=5.

Puzmanova:2006:BRC

[Puž06a] Rita Pužmanová. Book reviews: Cisco mobile IP configuration, troubleshooting, and management. IEEE Distributed Systems Online, 7(4):??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/04/o4006.pdf.

Puzmanova:2006:TFM

[Puz06b] Rita Puzmanová. Toward full mobility in converged networks. IEEE Distributed Systems Online, 7(8):??, August 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/08/
                                        o8004.pdf.

Puzmanova:2007:BRN

[Puz07a] Rita Puzmanová. Book reviews: Networks’ importance in Grids. IEEE Distributed Systems Online, 8(2):??, February 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/02/o2004.pdf.

Puzmanova:2007:GEN

[Puž07b] Rita Pužmanová. Grids enter a new generation. IEEE Distributed Systems Online, 8(7):??, July 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/07/
                                        mds2007070007.pdf.

Puzmanova:2007:VYN

[Puz07c] Rita Puzmanová. Voilà, your network is IPv6 enabled! IEEE Distributed Systems Online, 8(4):??, April 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/04/
o4006.pdf.

Qin:2001:FTS

[Qin01] Xiao Qin. Fault tolerant support in real-time collaborative editing systems. IEEE Distributed Systems Online, 2 (1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0101/departments/res0101_
print.htm.

Ramaswamy:2001:PAL

[Ram01] Ramkumar Ramaswamy. Probabilistic application-level connection scheduling in Web servers. IEEE Distributed Systems Online, 2(8), 2001. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://dsonline.computer.
org/0108/features/ram0108_
new.pdf.

Rao:2008:DSA

[Rao08] S. Rao. Distributed systems:

An algorithmic approach. IEEE Distributed Systems Online, 9 (11):??, November 2008. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

Roman:2001:RMY

[RCK01] Manuel Román, Roy H. Campbell, and Fabio Kon. Reflective middleware: From your desk to your hand. IEEE Distributed Systems Online, 2(5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/rom0105_
print.htm.

Rouvoy:2008:SAP

[RCS08] Romain Rouvoy, Denis Conan, and Lionel Seinturier. Software architecture patterns for a context-processing middleware framework. IEEE Distributed Systems Online, 9(6):??, June 2008. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2008/06/mds2008060001.pdf.

Ribeiro:2007:SPA

[RdCVM07] José Geraldo Ribeiro, Jr., Glauber Tadeu do Carmo, Marco Tílio Valente, and Nabor C. Mendonça. Smart proxies for accessing replicated Web services. IEEE Distributed Systems Online, 8(12): ??, December 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/12/
mds2007120001.pdf.

Regehr:2006:ERT

[Reg06] John Regehr. From the Editor: Real-time and embedded systems — teaching reliability. IEEE Distributed Systems Online, 7(5):??, May 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/05/o5002.pdf.

Rana:2004:TRD

[RH04a] O. F. Rana and A. Hinze. Trust and reputation in dynamic scientific communities. IEEE Distributed Systems Online, 5(1): ??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28452/01270714.pdf?isnumber=
28452&prod=JNL&arnumber=1270714&
arSt=&ared=&arAuthor=Rana%
2C+O.F.%3B+Hinze%2C+A.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
28452&arnumber=1270714&count=
8&index=3;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1270714.

Reiner:2004:OML

[RH04b] B. Reiner and K. Hahn. Optimized management of largescale data sets stored on tertiary storage systems. IEEE Distributed Systems Online, 5 (5):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28996/01306404.pdf?isnumber=

                                        28996&prod=JNL&arnumber=1306404&
                                        arSt=&ared=&arAuthor=Reiner%
                                        2C+B.%3B+Hahn%2C+K.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=28996&
                                        arnumber=1306404&count=5&
                                        index=0;   http://ieeexplore.
                                        ieee.org/xpls/references.
                                        jsp?arnumber=1306404.

Rimassa:2002:WWI

[Rim02]  Giovanni Rimassa.  Wired-

wireless integration: a middleware perspective. IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0209/d/w5icon.htm.

Rana:2006:BEM

[RK06] Omer F. Rana and Jeffrey O. Kephart. Building effective multivendor autonomic computing systems. IEEE Distributed Systems Online, 7(9): ??, September 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/09/
                                        o9003.pdf.

Rodriguez:2002:ISU

[Rod02] Martín González Rodríguez. IEEE software: Usability & navigability: Automatic datagathering agents for remote navigability testing. IEEE Distributed Systems Online, 3 (12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0212/f/s6_print.htm.

Rossi:2003:BRA

[Ros03] Bruno Rossi. Book review:

ASN.1’s best secrets: a review of ASN.1: Communication Between Heterogeneous Systems by Olivier Dubuisson. IEEE Distributed Systems Online, 4(2), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0302/d/bks_b.htm.

Rose:2006:BRE

[Ros06] Billy Rose. Book reviews: Extensible middleware. IEEE Distributed Systems Online, 7 (2):??, February 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/02/
o2007.pdf.

Ritter:2006:ISP

[RSL06] Tom Ritter, Rudolf Schreiner, and Ulrich Lang. Integrating security policies via container portable interceptors. IEEE Distributed Systems Online, 7 (7):??, July 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/07/o7001.pdf.

Sandhu:2002:IIC

[San02] Ravi Sandhu. IEEE Internet computing: Guest Editor’s introduction: The technology of trust. IEEE Distributed Systems Online, 3 (11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/f/w6gei.htm.

Sandhu:2003:IIC

[San03]  Ravi Sandhu.  IEEE In-

ternet computing: Security track: Good enough security. IEEE Distributed Systems Online, 4(1), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0301/f/wp1trak.htm.

Sastry:2002:EIR

[Sas02a]  Shiva Sastry.  An excel-

lent introduction: a review of Introduction to Distributed Algorithms by Gerard Tel. IEEE Distributed Systems Online, 3(6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0206/departments/bks_
a.htm.

Sastry:2002:FSR

[Sas02b] Shiva Sastry. A fair success: a review of Pervasive Computing Handbook by Uwe Hansmann, Lothar Merk, Martin S. Nicklous, and Thomas Stober. IEEE Distributed Systems Online, 3(9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/d/bks_c.htm.

Satyanarayanan:2002:CMU

[Sat02a] M. Satyanarayanan. A catalyst for mobile and ubiquitous computing. IEEE Distributed Systems Online, 3 (3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/departments/b1eic.
htm.

Satyanarayanan:2002:CIC

[Sat02b] M. Satyanarayanan. Challenges in implementing a contextaware system. IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/f/b3eic.htm.

Satyanarayanan:2003:IPC

[Sat03]  M. Satyanarayanan.  IEEE

pervasive computing: From the Editor in Chief: Privacy: The Achilles Heel of pervasive computing. IEEE Distributed Systems Online, 4 (3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0303/f/b1eic.htm.

Senouci:2007:PMS

[SBR^{+}07] Benaoumeur Senouci, Aimen Bouchhima, Frédéric Rousseau, Frédéric Petrot, and Ahmed Jerraya. Prototyping multiprocessor system-on-chip applications: a platform-based approach. IEEE Distributed Systems Online, 8(5):??, May 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2007/05/o5002.pdf.

Schreiner:2001:DPT

[Sch01a] Keri Schreiner. Distributed projects tackle protein mystery. IEEE Distributed Systems Online, 2(1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0101/features/fol0101_
print.htm.

Schreiner:2001:NJN

[Sch01b] Keri Schreiner. NASA’s JPL Nanorover outposts project. IEEE Distributed Systems Online, 2(3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0103/features/nan0103_
print.htm.

Schreiner:2001:PPI

[Sch01c] Keri Schreiner. Putting the profit in Internet-based DS. IEEE Distributed Systems Online, 2(5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/dco0105_
print.htm.

Sedighi:2006:BRG

[Sed06] Art Sedighi. Book reviews:

The Grid from the ground up. IEEE Distributed Systems Online, 7(11):??, November 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/11/oy004.pdf.

Semeczko:2002:BRS

[Sem02a] George Semeczko. Book review:

Securing XML: a review of XML Security by Blake Dournaee. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/bks_c.htm.

Semeczko:2002:IWT

[Sem02b] George Semeczko. An introduction to Web technologies: a review of Internet & World Wide Web: How to Program, 2nd edition. IEEE Distributed Systems Online, 3(4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/departments/bks_
c.htm.

Semeczko:2002:TBW

[Sem02c] George Semeczko. Tips on building Web pages: a review of HTML & Web Design: Tips & Techniques by Kris Jamsa, Konrad King, and Andy Anderson. IEEE Distributed Systems Online, 3(7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/departments/bks_
a.htm.

Semeczko:2003:BRJ

[Sem03a] George Semeczko. Book review: a J2EE reference for your library: a review of J2EE: The Complete Reference by Jim Keogh. IEEE Distributed Systems Online, 4 (3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0303/d/bks_c.htm.

Semeczko:2003:BRA

[Sem03b] George Semeczko. Book review: Advanced E-business: a review of E-Business Applications: Technologies for Tomorrow’s Solutions, Jorge Gasós and Klaus-Dieter Thoben, eds. IEEE Distributed Systems Online, 4(11), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0311/d/bks_c.htm.

Semeczko:2003:BRU

[Sem03c] George Semeczko. Book review: Unleashing the power of JSP: a review of JSP 2.0: The Complete Reference by Phil Hanna. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/0308/d/bks_b.htm.

Semeczko:2006:APP

[Sem06] George Semeczko. Applying the power of patterns to the Web. IEEE Distributed Systems Online, 7(1):??, January 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/01/o1006.pdf.

Semeczko:2007:CGR

[Sem07]  George Semeczko.  Com-

ing to grips with RSS feeds and standards. IEEE Distributed Systems Online, 8(9): ??, September 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/09/
mds2007090006.pdf.

Sacramento:2004:MMD

[SER^{+}04] V. Sacramento, M. Endler, H. K. Rubinsztejn, L. S. Lima, K. Goncalves, F. N. Nascimento, and G. A. Bueno. MoCA: a middleware for developing collaborative applications for mobile users. IEEE Distributed Systems Online, 5(10): 2, October 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/29772/01355907.pdf?isnumber=
29772&prod=JNL&arnumber=1355907&
arSt=+2&ared=+2&arAuthor=+
Sacramento%2C+V.%3B++Endler%
2C+M.%3B++Rubinsztejn%2C+
H.K.%3B++Lima%2C+L.S.%3B++
Goncalves%2C+K.%3B++Nascimento%

                                        2C+F.N.%3B++Bueno%2C+G.A.;
                                         http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        29772&arnumber=1355907&count=
                                        3&index=1.

Skjelsvik:2004:DEN

[SGP04] K. S. Skjelsvik, V. Goebel, and T. Plagemann. Distributed event notification for mobile ad hoc networks. IEEE Distributed Systems Online, 5(8): 2, August 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8968/29550/01341360.pdf?isnumber=
                                        29550&prod=JNL&arnumber=1341360&
                                        arSt=+2&ared=+2&arAuthor=+
                                        Skjelsvik%2C+K.S.%3B++Goebel%
                                        2C+V.%3B++Plagemann%2C+T.;
                                         http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        29550&arnumber=1341360&count=
                                        5&index=1.

Shannon:2003:BIP

[Sha03] Meg Shannon. In brief: Interactive posters pump up advertising. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0310/d/brief.htm.

Shapiro:2006:DWW

[Sha06] Marc Shapiro. Distributed wisdom: What’s wrong with us? IEEE Distributed Systems Online, 7(5):??, May 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/05/
o5003.pdf.

Sheth:2001:RTK

[She01] Vikram Sheth. A reference tool for the knowledgeable XML user. IEEE Distributed Systems Online, 2(6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0106/departments/bks0106b.
htm.

Sheth:2002:GRB

[She02a] Purvi Sheth. A great reference book for networking professionals: a review of Internetworking with TCP/IP Principles, Protocols, and Architectures, 4th edition by Douglas E. Comer. IEEE Distributed Systems Online, 3(6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0206/departments/bks_
b.htm.

Sheth:2002:UIW

[She02b] Vikram Sheth. A useful introduction to WAP: a review of Practical WAP — Developing Applications for Wireless Web by Chris Bennett. IEEE Distributed Systems Online, 3(7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/departments/bks_
b.htm.

Shif:2001:DFI

[Shi01] Gil Alexander Shif. DAML: The foundation of an intelligent Web. IEEE Distributed Systems Online, 2(1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0101/features/dam0101.
                                        htm.

Shneiderman:2002:IMV

[Shn02] Ben Shneiderman. IEEE Multimedia: Visions and views: Meeting human needs with new digital imaging technologies. IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0211/d/up4vav.htm.

Stabell-Kulo:2006:ESC

[SK06] Tage Stabell-Kulø. From the Editor: Security community — blurring the line between authentication and identification. IEEE Distributed Systems Online, 7(2):??, February 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/02/o2002.pdf.

Stainforth:2002:DCP

[SKA^{+}02] Dave Stainforth, Jamie Kettleborough, Myles Allen, Mat Collins, Andy Heaps, and James Murphy. Distributed computing for public-interest climate modeling research.

IEEE Distributed Systems Online, 3(4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/departments/sci_
print.htm.

Singh:2002:MRT

[SMN^{+}02] Gurdip Singh, Masaaki Mizuno, Mitch Neilsen, Donald Lenhert, and Naiqian Zhang. A multidisciplinary real-time embeddedsystems curriculum. IEEE Distributed Systems Online, 3 (5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0205/departments/edu_
print.htm.

Smith:2002:GSP

[SOB02] Garry Smith, Hong Ong, and Mark Baker. A good starting point for XML: a review of XML: A Beginner’s Guide. IEEE Distributed Systems Online, 3(3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/departments/bks_
c.htm.

Srinivas:2006:DWD

[SRJ06] A. Vijay Srinivas, M. Venkateswara Reddy, and D. Janakiram. Distributed wisdom: Designing a replication service for large peer-to-peer data grids. IEEE Distributed Systems Online, 7 (3):??, March 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/03/
                                        o3002.pdf.

Shen:2006:ECC

[SS06] Haifeng Shen and Chengzheng Sun. From the Editors: Collaborative computing community

— leveraging single-user applications for multiuser distributed collaboration. IEEE Distributed Systems Online, 7(4):??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/04/o4002.pdf.

Stewart:2004:PDC

[SSD^{+}04] C. Stewart, Kai Shen, S. Dwarkadas, M. L. Scott, and Jian Yin. Profile-driven component placement for cluster-based online services. IEEE Distributed Systems Online, 5(10):1, October 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/29772/01355906.pdf?isnumber=
                                        29772&prod=JNL&arnumber=1355906&
                                        arSt=+1&ared=+1&arAuthor=+
                                        Stewart%2C+C.%3B++Kai+Shen%
                                        3B++Dwarkadas%2C+S.%3B++Scott%
                                        2C+M.L.%3B++Jian+Yin;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29772&
                                        arnumber=1355906&count=3&
                                        index=0.

Stoop:2007:GSD

[ST07] Swen Stoop and Andy Turner. Games and simulation: DS online’s newest topic area. IEEE Distributed Systems Online, 8 (10):??, October 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/10/
mds2007100001.pdf.

Stachurski:2002:PCR

[Sta02a]  Andrzej Stachurski.  Paral-

lel computing reaches to industry: a review of Industrial Strength Parallel Computing edited by Alice E. Koniges. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0205/departments/bks_
b.htm.

Stanford:2002:UPC

[Sta02b] Vince Stanford. Using pervasive computing to deliver elder care. IEEE Distributed Systems Online, 3(3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/departments/bp1app.
htm.

Stanford:2003:IPCb

[Sta03a] Vince Stanford. IEEE pervasive computing: Applications: Pervasive computing goes the last hundred feet with RFID systems. IEEE Distributed Systems Online, 4(6), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0306/d/b2app.htm.

Stanford:2003:IPCa

[Sta03b] Vince Stanford. IEEE pervasive computing: Applications: Pervasive computing puts food on the table. IEEE Distributed Systems Online, 4(3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0303/d/p1app_print.htm.

Schilit:2002:CWW

[STHK02] Bill N. Schilit, Jonathan Trevor, David M. Hilbert, and Tzu Khiau Koh. Computer: Wireless: Web interaction using very small Internet devices. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0210/f/sch_print.htm.

Stirling:2003:BRT

[Sti03] Dale Stirling. Book review: a thorough look at Internet robots: a review of Beyond Webcams: An Introduction to Online Robots, Ken Goldberg and Roland Siegwart, eds. IEEE Distributed Systems Online, 4(3), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0303/d/bks_a.htm.

Stone:2001:B

[Sto01]  Adam Stone.  Bluetooth.

IEEE Distributed Systems Online, 2(7), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0107/departments/blu0107_
print.htm.

Stone:2002:BBB

[Sto02a] Adam Stone. In brief: Building better storage, one brick at a time. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/brief1.htm.

Stone:2002:BGG

[Sto02b] Adam Stone. In brief: Gaming on the Grid. IEEE Distributed Systems Online, 3 (11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/d/brief2.htm.

Stone:2002:BGB

[Sto02c] Adam Stone. In brief: Getting Bluetooth back on the road. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/brief2.htm.

Stone:2002:BMW

[Sto02d] Adam Stone. In brief: Majoring in wireless. IEEE Distributed Systems Online, 3 (11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/d/brief1.htm.

Stone:2002:NPA

[Sto02e] Adam Stone. .Net: a problem or an answer for opensourcing demands? IEEE Distributed Systems Online, 3 (3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/departments/new.
htm.

Stone:2002:NWS

[Sto02f] Adam Stone. The new Wireless Standard. IEEE Distributed Systems Online, 3 (8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0208/f/news_print.htm.

Stone:2002:NWW

[Sto02g] Adam Stone. News: Weaving the wi-fi blanket. IEEE Distributed Systems Online, 3 (10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/f/news.htm.

Stone:2002:OSW

[Sto02h]  Adam Stone.  OMA to

safeguard wireless content. IEEE Distributed Systems Online, 3(9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/d/brief2.htm.

Stone:2002:PPL

[Sto02i] Adam Stone. Pentagon policies limit wireless use. IEEE Distributed Systems Online, 3 (9), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0209/d/brief1.htm.

Stone:2002:WE

[Sto02j] Adam Stone. Wireless education. IEEE Distributed Systems Online, 3(1), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0201/departments/new_
1.htm.

Stone:2003:BCS

[Sto03a] Adam Stone. In brief: Canning spam. IEEE Distributed Systems Online, 4 (4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/d/brief.htm.

Stone:2003:BCW

[Sto03b] Adam Stone. In brief: Curbing war drivers. IEEE Distributed Systems Online, 4 (7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/d/brief.htm.

Stone:2003:NII

[Sto03c] Adam Stone. News: Internet invasion—is the Web secure? IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/f/News_print.htm.

Stone:2004:IWN

[Sto04] A. Stone. Investigating wireless networks with WHYNET. IEEE Distributed Systems Online, 5(4):??, April 2004. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/iel5/8968/28913/01301254.
                                        pdf?isnumber=28913&prod=JNL&
                                        arnumber=1301254&arSt=&ared=
                                        &arAuthor=Stone%2C+A.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=28913&
                                        arnumber=1301254&count=5&
                                        index=1.

Stockinger:2006:GCC

[Sto06a] Heinz Stockinger. Grid computing: a critical discussion on business applicability. IEEE Distributed Systems Online, 7 (6):??, June 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/06/o6002.pdf.

Stojanovic:2006:BRW

[Sto06b] Dragan Stojanovic. Book reviews: Web service-oriented project development from different perspectives. IEEE Distributed Systems Online, 7 (2):??, February 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/02/
o2008.pdf.

Stpiczynski:2002:PPO

[Stp02] Przemyslaw Stpiczynski. Parallel Programming in OpenMP helps novices: a review of Parallel Programming in OpenMP by Rohit Chandra, Leonardo Dagum, Dave Kohr, Dror Maydan, Jeff McDonald, and Ramesh Menon. IEEE Distributed Systems Online, 3 (8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0208/d/bks_a.htm.

Stpiczynski:2004:VRA

[Stp04] P. Stpiczynski. Valuable resources on aspects of scalable computing. IEEE Distributed Systems Online, 5(12):4, December 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/30113/01381278.pdf?isnumber=
30113&prod=JNL&arnumber=1381278&
arSt=+4&ared=+4&arAuthor=+
Stpiczynski%2C+P.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30113&
arnumber=1381278&count=4&
index=3.

Schelfthout:2006:MPB

[SWH06] Kurt Schelfthout, Danny Weyns, and Tom Holvoet. Middleware for protocol-based coordination in mobile applications. IEEE Distributed Systems Online, 7(8):??, August 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/08/
                                        o8001.pdf.

Shen:2007:CWC

[SYS07] Haifeng Shen, Zhonghua Yang, and Chengzheng Sun. Collaborative Web computing: From desktops to Webtops. IEEE Distributed Systems Online, 8(4):??, April 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2007/04/
                                        o4003.pdf.

Tanin:2007:TWS

[Tan07] Egemen Tanin. Teaching wireless sensor networks at the University of Melbourne. IEEE Distributed Systems Online, 8 (6):??, June 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2007/06/o6003.pdf.

Teller:2003:ICG

[TCB03] Seth Teller, Jiawen Chen, and Hari Balakrishnan. IEEE computer graphics & applications: Projects in VR: Pervasive pose-aware applications and infrastructure. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/f/gp4proj.htm.

Tohma:2004:IFT

[Toh04]  Y. Tohma.  Incorporat-

ing fault tolerance into an autonomic-computing environment. IEEE Distributed Systems Online, 5(2):3/1–3/12, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
8968/28454/01270783.pdf?isnumber=
28454&prod=JNL&arnumber=1270783&
arSt=+3%2F1&ared=+3%2F12&
arAuthor=+Tohma%2C+Y.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270783&count=6&
index=1;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1270783.

Touch:2002:PN

[Tou02]  Joseph D. Touch.  Those

pesky NATs. IEEE Distributed Systems Online, 3 (7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/departments/wp4icon.
htm.

Talia:2003:IIC

[TT03] Domenico Talia and Paolo Trunfio. IEEE Internet computing: Peer to peer toward a synergy between P2P and Grids.

IEEE Distributed Systems Online, 4(7), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0307/d/wp4p2p.htm.

Uguz:2002:LHR

[Ugu02] Agah Uguz. Learning HTML: a review of How to Do Everything with HTML by James H. Pence. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0205/departments/bks_
                                        a.htm.

Ulm:2004:WWW

[Ulm04] D. R. Ulm. World wide wisdom. IEEE Distributed Systems Online, 5(9):5, September 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/29771/01355905.pdf?isnumber=
                                        29771&prod=JNL&arnumber=1355905&
                                        arSt=+5&ared=+5&arAuthor=+
                                        Ulm%2C+D.R.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29771&arnumber=1355905&
                                        count=4&index=3.

Vargas:2006:DCE

[Var06] Luis Vargas. Database challenges in enterprise information sharing. IEEE Distributed Systems Online, 7(10):??, October 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.

computer.org/comp/mags/ds/
2006/10/ox002.pdf.

Veglis:2002:PCR

[Veg02a] Andreas Veglis. Programming with ColdFusion 5.0: a review of Programming Cold-Fusion by Rob Brooks-Bilson. IEEE Distributed Systems Online, 3(7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0207/departments/bks_
c.htm.

Veglis:2002:WBT

[Veg02b]  Andreas A. Veglis.  Web-

based teaching systems. IEEE Distributed Systems Online, 3 (4), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0204/departments/edu_
print.htm.

Veglis:2003:BRH

[Veg03]  Andreas Veglis.  Book re-

view: Harvesting the power of Google: a review of Google Hacks: 100 Industrial-Strength Tips & Tools by Tara Calishain and Rael Dornfes. IEEE Distributed Systems Online, 4(10), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0310/d/bks_c.htm.

Veglis:2006:OSW

[Veg06] Andreas Veglis. Open source Web software for Windows users. IEEE Distributed Systems Online, 7(7):??, July 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/ds/
2006/07/o7005.pdf.

Veglis:2007:LDA

[Veg07] Andreas Veglis. Learn to design attractive Web sites. IEEE Distributed Systems Online, 8 (11):??, November 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/11/
mds2007110006.pdf.

vanHarmelen:2004:SWW

[vH04] F. van Harmelen. The Semantic Web: what, why, how, and when. IEEE Distributed Systems Online, 5(3):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28687/01285880.pdf?isnumber=
28687&prod=JNL&arnumber=1285880&
arSt=&ared=&arAuthor=van+
Harmelen%2C+F.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28687&
arnumber=1285880&count=7&
index=3.

Villa:2002:CTI

[Vil02] Juan Pablo Villa. CDMA trends & implementations: a review of IS-95 CDMA & cdma2000: Cellular/PCS Systems Implementation by Vijay K. Garg.

IEEE Distributed Systems Online, 3(8), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0208/d/bks_c.htm.

Villa:2003:BRD

[Vil03] Juan Pablo Villa. Book review: Dial 911 for Ross: a review of Security Engineering by Ross J. Anderson. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/d/bks_b.htm.

Vinoski:2002:ICT

[Vin02] Steve Vinoski. Internet computing: Toward integration: Middleware “Dark Matter”. IEEE Distributed Systems Online, 3(10), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0210/d/wp5midd.htm.

Vinoski:2003:IICc

[Vin03a] Steve Vinoski. IEEE Internet computing: Toward integration: Invocation styles. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/d/w4midp.htm.

Vinoski:2003:IICb

[Vin03b] Steve Vinoski. IEEE Internet computing: Toward integration: It’s just a mapping problem. IEEE Distributed Systems Online, 4(5), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/d/wp3mid.htm.

Vinoski:2003:IICa

[Vin03c] Steve Vinoski. IEEE Internet computing: Toward integration: The performance presumption. IEEE Distributed Systems Online, 4(4), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0304/d/wp2midd.htm.

Vitenberg:2001:CSC

[Vit01] Roman Vitenberg. Caching support for CORBA objects. IEEE Distributed Systems Online, 2(3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0103/departments/res0103_
                                        print.htm.

Vollset:2002:BRP

[Vol02]  Einar Vollset.  Book re-

view: a peek into the future?: a review of The Intelligent Wireless Web by H. Peter Alesso and Craig F. Smith. IEEE Distributed Systems Online, 3(11), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0211/d/bks_b.htm.

Voth:2003:BWY

[Vot03]  Danna Voth.  In brief:

“Where Are You?” watches. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/0308/d/brief1.htm.

Verhanneman:2006:MAC

[VPW^{+}06] Tine Verhanneman, Frank Piessens, Bart De Win, Eddy Truyen, and Wouter Joosen. A modular access control service for supporting applicationspecific policies. IEEE Distributed Systems Online, 7(6): ??, June 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/06/o6001.pdf.

Vrenios:2004:PPC

[Vre04] A. Vrenios. Parallel Programming in C with MPI and OpenMP [book review]. IEEE Distributed Systems Online, 5 (1):7.1–7.3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.org/
iel5/8968/28452/01270716.
pdf?isnumber=28452&prod=JNL&
arnumber=1270716&arSt=+7.
1&ared=+7.3&arAuthor=Vrenios%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?

                                        isnumber=28452&arnumber=1270716&
                                        count=8&index=5.

vanSteen:2000:CS

[vS00]  Maarten van Steen.  Call

for submissions. IEEE Distributed Systems Online, 1 (2), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds200/ds2res.
                                        htm.

vanSteen:2001:DSB

[vS01]  Maarten van Steen.  Dis-

tributed systems book covers all the relevant issues. IEEE Distributed Systems Online, 2(1), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0101/departments/bks0101.
                                        htm.

vanSteen:2004:CSD

[vS04] M. van Steen. On the complexity of simple distributed systems. IEEE Distributed Systems Online, 5(5):??, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/28996/01306405.pdf?isnumber=
                                        28996&prod=JNL&arnumber=1306405&
                                        arSt=&ared=&arAuthor=van+
                                        Steen%2C+M.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=28996&arnumber=1306405&
                                        count=5&index=1.

Venkatasubramanian:2001:SFM

[VT01]  Nalini  Venkatasubramanian

and Carolyn L. Talcott. A semantic framework for modeling and reasoning about reflective middleware. IEEE Distributed Systems Online, 2 (6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0106/features/ven0106.
pdf.

Weatherford:2000:WGG

[Wea00] Margaret Weatherford. Where is Guide going? IEEE Distributed Systems Online, 1 (2), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds200/ds2updprint.
htm.

Weatherford:2001:AMS

[Wea01a] Margaret Weatherford. AIDE:

Making space for collaborative engineering. IEEE Distributed Systems Online, 2 (6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0106/departments/new0106.
htm.

Weatherford:2001:PAJ

[Wea01b] Margaret Weatherford. P2P acclaimed by a jury of peers. IEEE Distributed Systems Online, 2(3), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0103/features/p2p0103_
                                        print.htm.

Weatherford:2002:MF

[Wea02] Margaret Weatherford. Mining for fraud. IEEE Distributed Systems Online, 3 (7), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0207/departments/news_
                                        IS.htm.

Weissman:2000:GC

[Wei00] Jon B. Weissman. Grids in the classroom. IEEE Distributed Systems Online, 1 (3), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/archives/ds300/ds3eduprint.
                                        htm.

Wilkinson:2006:TGCa

[WF06a] Barry Wilkinson and Clayton Ferner. Teaching Grid computing in North Carolina: Part I. IEEE Distributed Systems Online, 7(6):??, June 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/06/
                                        o6003.pdf.

Wilkinson:2006:TGCb

[WF06b] Barry Wilkinson and Clayton Ferner. Teaching Grid computing in North Carolina: Part II.

IEEE Distributed Systems Online, 7(7):??, July 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2006/07/
o7003.pdf.

Willmott:2004:DSS

[Wil04] S. Willmott. On distributed systems and social engineering. IEEE Distributed Systems Online, 5(11):1, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/30052/01377088.pdf?isnumber=
30052&prod=JNL&arnumber=1377088&
arSt=+1&ared=+1&arAuthor=
+Willmott%2C+S.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30052&
arnumber=1377088&count=5&
index=0;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1377088.

Williams:2006:BRB

[Wil06] Godfried Williams. Book reviews: Breaking security design bottlenecks in payment technologies. IEEE Distributed Systems Online, 7(11):??, November 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//csdl.computer.org/comp/
mags/ds/2006/11/oy005.pdf.

Woodman:2002:BRI

[Woo02] Simon Woodman. Book review: An introduction to the world of Web services: a review of Web Services: Building Blocks for Distributed Applications by Graham Glass. IEEE Distributed Systems Online, 3(12), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0212/d/bks_b.htm.

Want:2000:UETa

[WR00a] Roy Want and Daniel M. Russell. Ubiquitous electronic tagging. IEEE Distributed Systems Online, 1(3), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds200/ds2wan.
htm.

Want:2000:UETb

[WR00b] Roy Want and Daniel M. Russell. Ubiquitous electronic tagging. IEEE Distributed Systems Online, 1(3), 2000. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/archives/ds200/ds2wan.
htm.

Wang:2001:ARM

[WSKP01] Nanbor Wang, Douglas C. Schmidt, Michael Kircher, and Kirthika Parameswaran. Adaptive and reflective middleware for QoS-enabled CCM applications. IEEE Distributed Systems Online, 2 (5), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0105/features/wan0105_
print.htm.

Xue:2001:NVA

[Xue01] Zhihui (Jeff) Xue. A narrow view of agents and conflict management. IEEE Distributed Systems Online, 2 (6), 2001. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0106/departments/bks0106a.
htm.

Xue:2002:DOS

[Xue02a] Zhihui (Jeff) Xue. Distributed operating systems: Classical and current techniques: a review of Distributed Operating Systems: Concepts and Practice. IEEE Distributed Systems Online, 3(3), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0203/departments/bks_
b.htm.

Xue:2002:IVA

[Xue02b] Zhihui (Jeff) Xue. An integrated view of algorithms: a review of Algorithms Sequential and Parallel: A Unified Approach by Russ Miller and Laurence Boxer. IEEE Distributed Systems Online, 3(5), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0205/departments/bks_
c.htm.

Yan:2005:EPC

[Yan05] Lu Yan. Enable pervasive computing with Java. IEEE Distributed Systems Online, 6(6): ??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2005/06/o6004.pdf.

Yan:2006:BRW

[Yan06] Lu Yan. Book reviews: Wireless sensor networks — past, present, and future. IEEE Distributed Systems Online, 7(5): ??, May 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/ds/2006/05/o5007.pdf.

Yan:2007:TUC

[Yan07] Lu Yan. On teaching ubiquitous computing. IEEE Distributed Systems Online, 8(7):??, July 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2007/07/mds2007070004.pdf.

Yehezkael:2006:FAO

[Yeh06] R. B. Yehezkael. Flexible algorithms: Overview of a beginners’ course. IEEE Distributed Systems Online, 7(11): ??, November 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/ds/2006/11/
                                        oy002.pdf.

Yehezkael:2007:FAF

[Yeh07]  R. B. Yehezkael.  Flexible

algorithms: Fragments from a beginners’ course. IEEE Distributed Systems Online, 8 (2):??, February 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/02/
o2001.pdf.

Yang:2004:GEI

[YPSW04] L. T. Yang, M. Paprzycki, Xiaohui Shen, and Xingfu Wu. Guest editors’ introduction [data-intensive computing]. IEEE Distributed Systems Online, 5(1):1.1–1.3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28452/01270710.pdf?isnumber=
28452&prod=JNL&arnumber=1270710&
arSt=+1.1&ared=+1.3&arAuthor=
Yang%2C+L.T.%3B+Paprzycki%
2C+M.%3B+Xiaohui+Shen%3B+
Xingfu+Wu;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=28452&arnumber=1270710&
count=8&index=0.

Zalech:2003:BRW

[Zal03] Wojciech Zalech. Book review: What hides behind the Internet-HTTP protocol. IEEE Distributed Systems Online, 4(8), 2003. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0308/0306/d/bks_a.htm.

Zalech:2004:EAN

[Zal04] Z. Zalech. Essential ASP.NET2. IEEE Distributed Systems Online, 5(12):3, December 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/iel5/8968/30113/01381277.
                                        pdf?isnumber=30113&prod=JNL&
                                        arnumber=1381277&arSt=+3&
                                        ared=+3&arAuthor=+Zalech%
                                        2C+Z.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30113&arnumber=1381277&
                                        count=4&index=2.

ZenilC:2002:GJP

[Zen02]  Héctor Zenil C.  Good

JavaServer pages introduction: a review of Web Development with JavaServer Pages, 2nd edition by Duane K. Fields and Mark A. Kolb. IEEE Distributed Systems Online, 3(6), 2002. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0206/departments/bks_
                                        c.htm.

Zheng:2006:SRS

[ZN06] Pei Zheng and Lionel M. Ni. Spotlight: The rise of the smart phone. IEEE Distributed Systems Online, 7(3):??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/ds/
                                        2006/03/o3003.pdf.